### v1

# Learning from Imperfect Human Feedback: a Tale from Corruption-Robust $\text{Dueling}^{*}$

Abstract
This paper studies Learning from Imperfect Human Feedback (LIHF), motivated by humans’ potential irrationality or imperfect perception of true preference. We revisit the classic dueling bandit problem as a model of learning from comparative human feedback, and enrich it by casting the imperfection in human feedback as agnostic corruption to user utilities. We start by identifying the fundamental limits of LIHF and prove a regret lower bound of $\Omega(\max\{T^{1/2},C\})$, even when the total corruption $C$ is known and when the corruption decays gracefully over time (i.e., user feedback becomes increasingly more accurate). We then turn to design robust algorithms applicable in real-world scenarios with arbitrary corruption and unknown $C$. Our key finding is that gradient-based algorithms enjoy a smooth efficiency-robustness tradeoff under corruption by varying their learning rates. Specifically, under general concave user utility, Dueling Bandit Gradient Descent (DBGD) of Yue and Joachims (2009) can be tuned to achieve regret $O(T^{1-\alpha}+T^{\alpha}C)$ for any given parameter $\alpha\in(0,\frac{1}{4}]$. Additionally, this result enables us to pin down the regret lower bound of the standard DBGD (the $\alpha=1/4$ case) as $\Omega(T^{3/4})$ for the first time, to the best of our knowledge. For strongly concave user utility we show a better tradeoff: there is an algorithm that achieves $O(T^{\alpha}+T^{\frac{1}{2}(1-\alpha)}C)$ for any given $\alpha\in[\frac{1}{2},1)$. Our theoretical insights are corroborated by extensive experiments on real-world recommendation data.
1 Introduction
Many real-world problems, such as personalized recommendation (Yue and Joachims, 2009; Immorlica et al., 2020; Yao et al., 2022a) and fine-tuning generative models (Bai et al., 2022; Casper et al., 2023; Han et al., 2024), require learning from human feedback. Expressing preferences as numerical values or concrete functions is generally challenging for humans. Therefore, an approach that has achieved significant real-world success is to learn humans’ preferences by eliciting their comparative feedback between two options (Bai et al., 2022). A natural theoretical framework capturing such learning task is the seminal dueling bandits framework introduced by Yue and Joachims (2009). The dueling bandit problem features an sequential online learning problem, during which a pair of actions are selected at each round and only their comparison result is revealed to the learner. The comparison outcome between two actions is modeled using a utility function of actions, alongside a link function that determines the probability of each action winning based on their utility difference. This modeling approach is termed as utility-based dueling bandit and achieves great success in generating summaries closely aligned with human preferences (Stiennon et al., 2020).
However, human feedback is often imperfect, a crucial factor that has been largely overlooked in previous studies of dueling bandit. Ample behavioral studies show that humans often refine their preferences through interactions with the system, navigating a complex tradeoff between exploration and exploitation (Wilson et al., 2014; Cohen et al., 2007). Furthermore, their response errors may exhibit systematic patterns rather than merely random noise. As we are all aware of, humans are not always rational (Posner, 1997) neither perfectly know our preferences (Pu et al., 2012).
In this paper, we study learning from imperfect human feedback (LIHF) by casting the imperfection in the feedback as certain form of corruption to the utilities. This leads us to study a natural variant of the continuous dueling bandit with corrupted utilities. Notably, our study differs from recent robust algorithms designed for corrupted dueling bandits (Di et al., 2024; Saha and Gaillard, 2022; Komiyama et al., 2015) in three key aspects. Firstly, these previous algorithms all focus on a finite action space whereas, to the best of our knowledge, ours is the first to investigate a robust dueling bandit algorithm under corruption for continuous action space. This focus is motivated by the fact that many real-world applications have continuous or extremely dense action spaces, such as token embeddings or content recommendations (Chen et al., 2021, 2019). Second, our algorithms are designed for agnostic corruption without knowing the corruption levels, whereas Di et al. (2024) requires knowledge of the total corruption amount. This focus on agnostic corruption is motivated by realistic considerations, because systems often cannot accurately discern the irrational or imperfect nature of human feedback. Third, our corruption models are very different from the model studied by Di et al. (2024); Saha and Gaillard (2022); Komiyama et al. (2015). The corruption in their model contaminates feedback by flipping the realized outcomes of the duels, whereas the corruption in our model is on utilities which consequently only alter the winning probabilities of the duels. This difference is due to our different motivation of corruption, which is to capture imperfect human behavior while not intentional adversarial attack (see more illustration in Section 3).
Our Contributions. By modeling LIHF as a continuous dueling bandit under agnostic corruption, we start by characterizing its intrinsic hardness. We prove a regret lower bound of $\Omega(d\max\{\sqrt{T},C\})$, even when the total corruption $C$ is known and when corruption decays gradually over time. Additionally, we prove that this lower bound result is tight by exhibiting an algorithm that indeed achieves a matching upper bound for the aforementioned setting. Next, we turn to design robust algorithms that are applicable to arbitrary and agnostic corruption with unknown $C$. Our key finding is that gradient-based algorithms offer a smooth efficiency-robustness tradeoff under corruption by carefully tuning their learning rates. Specifically, let $d$ denote the problem dimension. For any given $\alpha\in(0,1/4]$, Dueling Bandit Gradient Descent (DBGD) (Yue and Joachims, 2009) with carefully chosen learning rate (that depends on $\alpha$) can achieve a regret of $O(\sqrt{d}T^{1-\alpha}+\sqrt{d}T^{\alpha}C)$ under arbitrary concave user utility. That is, by decreasing $\alpha$ from $1/4$ to 00, the regret under no corruption (i.e., learning efficiency) decays from $\sqrt{d}T^{3/4}$ to $\sqrt{d}T$ but the algorithm’s tolerance to total amount of corruption increases from $O(T^{3/4})$ to $O(T)$ (i.e., more robustness). Perhaps surprisingly, this result turns out to allow us to pin down a regret lower bound for the standard version of DBGD (the $\alpha=1/4$ case) as $\Omega(T^{3/4})$, marking the first tight lower bound result for the DBGD algorithm to the best of our knowledge. For strongly concave user utility, we show that an improved tradeoff is possible: Noisy Comparison based Stochastic Mirror Descent (NC-SMD) (Kumagai, 2017) can be refined to achieve a regret of $O(dT^{\alpha}+\sqrt{d}T^{\frac{1-\alpha}{2}}C+dC)$ for any given $\alpha\in[1/2,1)$. We remark that similar robustness-accuracy tradeoff has been studied classification problem with adversarial examples, both empirically (Tsipras et al., 2018) and theoretically (Zhang et al., 2019). However, to the best of our knowledge, this is the first time such tradeoff is formally analyzed for online learning. Finally, we conduct extensive experiments on real-world recommendation data to validate our theoretical insights.
2 Related Work
Hardness of Learning from Imperfect Human Feedback: To design algorithms robust to adversarial corruption, it is crucial to understand the intrinsic hardness of learning from corrupted feedback. Lattimore and Szepesvári (2020) demonstrates that the minimax regret lower bound for stochastic linear bandits is $\Omega(d\sqrt{T})$. Utilizing the regret lower bound for bandit convex optimization (Shamir, 2013), Kumagai (2017) illustrates that dueling bandits with a strongly concave utility function achieve a regret lower bound of $\Omega(d\sqrt{T})$. In scenarios of adversarial corruption, Bogunovic et al. (2020) proves that any algorithm faces regret at least $\Omega(dC)$ for stochastic linear bandits. However, currently there is no literature that has explored the intrinsic hardness of learning from corruption which decays gradually over time (Wilson et al., 2014; Kumagai, 2017; Cohen et al., 2007), despite its prevalence in learning from imperfect human feedback. It still remains unknown whether learning from such corruption is fundamentally easier than learning from arbitrary corruption.
Stochastic Bandits with Unknown Corruption: Applying the principle of optimism in face of uncertainty and constructing estimators involving uncertainty weighting, algorithms devised by He et al. (2022) and Ye et al. (2024a) attain nearly optimal regret bounds for both linear and nonlinear contextual bandits in presence of adversarial corruption. By leveraging the concept of phase elimination, Bogunovic et al. (2022) introduces a robust algorithm capable of handling corruption in rewards drawn from Gaussian processes. However, the designs of these algorithms require explicit knowledge of the total corruption level $C$ as an input. Gupta et al. (2019) employ novel sampling strategy to channel resources to arms with better performance and craft algorithms robust to agnostic corruption, yet restricted to finite action sets. In scenarios involving continuous action set with unknown corruption level, even though algorithms designed by Ding et al. (2022) and Bogunovic et al. (2020) have sublinear regret guarantee, the regret is at least $C$ times worse than the regret without corruption. It remains as an open problem of whether better regret guarantee is achievable—that is, affording unknown corruption level more than $\Omega(\sqrt{T})$ while maintaining $O(\sqrt{T})$ regret guarantee in scenario without corruption.
Continuous Dueling Bandits with Unknown Corruption: In addition to stochastic bandits, there exists an extensive body of literature studying the dueling bandit problem with unknown corruption, given its prevalence in real-world application (Bengs et al., 2021b; Sui et al., 2018). The literature on dueling bandits with adversarial corruption has been initially explored by Agarwal et al. (2021) and subsequently enhanced by Saha and Gaillard (2022), which achieves optimality in term of $C$. However, their work is solely applicable to finite action sets, which may not be feasible in many real-world scenarios that have continuous or extremely dense action spaces, such as token embeddings or content recommendations (Chen et al., 2021, 2019). Ailon et al. (2014) introduced a reduction method that translates existing results from conventional Multi-Armed Bandit algorithms to Dueling Bandits. This reduction method is applicable to both finite and continuous action spaces, albeit requiring a linear link function. DBGD (Yue and Joachims, 2009) and NC-SMD (Kumagai, 2017) enjoy flexibility in choosing the link function but only work for continuous dueling bandits. It still remains as an open problem for developing provably efficient algorithm capable of learning from corrupted dueling feedback with unknown corruption in continuous action space. This serves as the driving force behind our theoretical analysis.
3 The Problem of Learning from Imperfect Human Feedback (LIHF)
Conventional Notation. Throughout the paper, for a positive integer $T$, we use $[T]$ to denote $\{1,2,\ldots,T\}$. We use $\mathcal{P}_{\mathcal{A}}(a)$ to denote the projection of action $a$ on $\mathcal{A}$. We use standard asymptotic notations including $O(\cdot)$, $\Omega(\cdot)$, $\Theta(\cdot)$. We use $\tilde{O}(\cdot)$, $\tilde{\Omega}(\cdot)$, $\tilde{\Theta}(\cdot)$ to hide logarithmic factors. We use $\|\cdot\|_{\infty}$ to define infinity norm. $\text{Reg}_{T}$ represents regret with dueling feedback over $T$ rounds.
To capture learning from imperfect (comparative) human feedback, we now introduce a natural variant of the seminal dueling bandit framework (Yue and Joachims, 2009) by integrating the “imperfections” in human feedback as certain corruption to human agents’ feedback. An immediate critique one might have is that arbitrarily adversarial corruption seems too pessimistic as a way to model humans’ inaccuracies which are often caused by uncertainty or ignorance, rather than targeted attacks. While this may appear true at the first glance, we provide strong theoretical and empirical evidences that corruption caused by uncertainties generally is as hard as arbitrary corruption. For such reasons, we coin the following problem Learning from Imperfect Human Feedback (LIHF).
Basics of Dueling Bandits and Assumptions. We consider the standard dueling bandit framework with continuous action set $\mathcal{A}\subset\mathbb{R}^{d}$ (Yue and Joachims, 2009). At each round $t\in[T]$, the learner (henceforth denoted as $\mathcal{L}$) chooses two actions $a_{t},a^{\prime}_{t}$ to present to some human agent, henceforth denoted as the “user”. The user receives utility $\mu(a_{t}),\mu(a^{\prime}_{t})$ from the two actions respectively, and will pick one of these actions following a link function $\sigma(\cdot)$. That is, in absence of any corruption, the user picks action $a_{t}$ with probability $\sigma(\mu(a_{t})-\mu(a^{\prime}_{t}))$, and action $a^{\prime}_{t}$ otherwise. As clear from its definition, this choice probability depends only on the utility difference $\mu(a_{t})-\mu(a^{\prime}_{t})$. Notably, user’s randomized action does carry information about underlying utility function $\mu$, which is why the action is also called the “human feedback” (Bengs et al., 2021a; Maystre and Grossglauser, 2017; Ailon et al., 2014). Following standard assumptions in this literature (Kumagai, 2017; Yue and Joachims, 2009), we also assume
  1. 1.
    the continuous action space $\mathcal{A}$ contains the origin, is compact, convex and is contained in a $d$-dimensional ball of radius $R$;
  2. 2.
    the utility function $\mu:\mathcal{A}\rightarrow\mathbb{R}$ is concave, differentiable, $L_{\mu}$-Lipschitz and admits unique optimal within set $\mathcal{A}$, denoted as $a^{*}:=\arg\max_{a\in\mathcal{A}}\mu(a)$;
  3. 3.
    the link function $\sigma:\mathbb{R}\rightarrow[0,1]$ is rotational-symmetric, $L_{\sigma}$-Lipschitz and its first derivative $\sigma^{\prime}$ is $L_{2}$-Lipschitz.
For convenience, we further let $L=L_{\sigma}L_{\mu}$. Notably, some previous works restrict link function to specifically be the logistic link function (Di et al., 2024; Xu et al., 2024; Saha, 2021). However, our results shall be applicable to any link functions that satisfy the natural assumptions above.
Inaccuracies in Human Feedback as Agnostic Corruption. Extending the standard dueling bandit framework above, we introduce a corruption term $c_{t}(a_{t},a^{\prime}_{t})$ to the utility difference in order to model potential inaccuracy in the user’s perception of underlying true utilities. For all our positive results, we always work with the situation where this corruption is agnostic in the sense that the learner does not know the value of each $c_{t}$ neither their accumulated sum. Such agnostic corruption is natural because in reality it is very difficult to truly estimate how imperfect a user is, hence we aim to develop algorithms whose execution is agnostic to such parameters. Let $a\succ a^{\prime}$ denote the event that the user chooses $a$ over $a^{\prime}$. Hence in our LIHF problem, the probability of event $a\succ a^{\prime}$ in presence of corruption $c(a,a^{\prime})$, denoted as ^$\hat{\mathbb{P}}(a\succ a^{\prime})$, can be expressed as follows ^$\hat{\mathbb{P}}(a\succ a^{\prime})=\sigma(\mu(a)-\mu(a^{\prime})+c(a,a^{\prime}))$. We further let ^$\hat{\mathcal{F}}(a_{t},a^{\prime}_{t})$ denote user’s preferential feedback under corruption, coined corrupted dueling feedback. Mathematically, ^$\hat{\mathcal{F}}(a_{t},a^{\prime}_{t})$ follows a binomial distribution with mean ^$\hat{\mathbb{P}}(a_{t}\succ a^{\prime}_{t})$, expressed by the following equation.
^$\hat{\mathcal{F}}(a_{t},a^{\prime}_{t}):=\begin{cases}1\hskip 2.84526pt\text{w.p.}\hskip 2.84526pt\hat{\mathbb{P}}(a_{t}\succ a^{\prime}_{t})\\ 0\hskip 2.84526pt\text{w.p.}\hskip 2.84526pt1-\hat{\mathbb{P}}(a_{t}\succ a^{\prime}_{t})\end{cases}$
The “hat” notation in $\hat{\mathcal{F}}$ and $\hat{\mathbb{P}}$ is to denote corruption. When there is no corruption, we use $\mathbb{P}(a\succ a^{\prime})=\sigma(\mu(a)-\mu(a^{\prime}))$ to denote the probability of event $a\succ a^{\prime}$ and $\mathcal{F}(a,a^{\prime})$ to denote corresponding dueling feedback. It is important to recognize the different roles of the link function $\sigma$ and utility corruption $c_{t}$$\sigma$ captures the inherent (often inevitable) randomness in human choice/behavior even given true utilities (Fishburn et al., 1979; Becker et al., 1963; Bradley and Terry, 1952), whereas our $c_{t}$ is introduced to model inaccurate perception of utilities due to not perfectly knowing her preferences and, importantly, can be alleviated or even eliminated once the user has tried sufficient actions.
Imperfect user feedback generally is not as wild as arbitrary adversarial corruption. A particularly natural pattern of human’s imperfection is that the inaccuracy in our estimations often diminishes as we try more and more actions. That is, humans learn along the way, though how fast the inaccuracy decreases often depends on our learning rate. This hence motivates us to consider the following notion of $\rho$-imperfect user which significantly constrain the corruption sequence $c_{1},c_{2},\cdots$, and force it to diminish at the rate $\rho$.
Definition 1 ($\rho$-Imperfect User).
The user feedback is said to be $\rho$-imperfect for some $\rho\in[0,1]$ if there exists a constant $C_{\kappa}>0$ such that the corruption $c_{t}$ satisfies $|c_{t}|\leq C_{\kappa}t^{\rho-1},\forall t\in[T]$. [^1] All our results naturally applies to $\rho<0$, which is a significantly easier situation since accumulated corruption is a constant in that case. We hence will not explicitly consider it in this paper.
For a $\rho$-imperfect user, the accumulated corruption $\sum_{\tau=1}^{t-1}|c_{\tau}|\leq C_{\kappa}t^{\rho}$ for every $t$, which is why this is coined $\rho$-imperfect. The diminishing inaccuracy of user estimation in Definition 1 is a natural model for user’s learning behavior and has already been used in various previous works to model user’s imperfect feedback (Wang et al., 2023; Yao et al., 2022a; Immorlica et al., 2020) (Details are presented at Appendix A). Note that agnostic corruption in this situation means the learner know $|c_{t}|\leq C_{\kappa}t^{\rho-1}$ for some $\rho$, but she does not know the value of $\rho$.
On the other hand, the corruption are said to be arbitrary up to certain budget $C$, if $\sum_{t=1}^{T}|c_{t}|\leq C$ for a given (hence fixed) time horizon $T$. Our designed (positive) learning algorithmic will have guaranteed regret upper bound under arbitrary corruption, though our lower bound (i.e., negative result) will hold even in the weaker $\rho$-imperfect user feedback.
We remark that another line of recent research studies a different type of corruption that directly flip the outcomes of the duels; i.e., flipping the realized event $a\succ a^{\prime}$ to its opposite $a^{\prime}\succ a$ (Di et al., 2024; Saha and Gaillard, 2022; Komiyama et al., 2015). These corruption models are motivated by the existence of true adversaries, e.g., malicious users in recommendation systems (Lykouris et al., 2018) or fraudulent clicks (Deshpande and Montanari, 2013), who have the capability of altering user choices. In contrast, only utilities are corrupted in our model. These two models are not comparable. For instance, the adversary can induce a deterministic event $a\succ a^{\prime}$ at any given round with one corruption, whereas the outcomes in our model is almost always probabilistic (unless the adversary uses an infinite budget $c_{t}$). Such modeling difference inherently comes from the different motivations of our LIHF problem. Specifically, our model is to capture imperfect human feedback due to inaccurate perception of utilities or ignorance of true preferences. For example, when recommender systems learn to find the optimal recommendation for a user, the user may not know her true preference at the beginning and often gradually refines her preferences through experiencing different items (Yao et al., 2022b). Similarly, when generative models like ChatGPT learn to generate personalized content for a user, the user may go through a similar procedure of preference refinement during interactions with the model. These modelling differences also necessitate our fundamentally different techniques. Our approaches are based on gradient-based methods whereas Di et al. (2024) and Saha (2021) employ the principle of optimistic in face of uncertainty and construct the confidence set of preferences through querying the maximum information pair. Their approaches inherently require the knowledge of total corruption level $C$. In contrast, the analysis of our gradient-based methods controls the impact of gradient bias (due to corruption) on regret. By adjusting its learning rate, our algorithm will gracefully “trade” its learning efficiency for more robustness against agnostic corruption. Moreover, our algorithm is suitable for infinite action space, whereas previous works of Di et al. (2024); Saha (2021) are for finite action spaces.
Learning goal: agnostic corruption-robust regret minimization. The goal of the learner is to optimize her sequential actions to minimize the following dueling regret against an arbitrary sequence of corruption $\{c_{t}\}_{t=1}^{T}$:
$\text{Reg}_{T}=\mathbb{E}\left(\sum^{T}_{t=1}\sigma(\mu(a^{*})-\mu(a_{t}))+\sigma(\mu(a^{*})-\mu(a^{\prime}_{t}))-1\right).$ (1)
This regret measure is also studied in many prior literature (Saha and Gaillard, 2022; Kumagai, 2017; Komiyama et al., 2015; Yue and Joachims, 2009). Importantly, all our algorithms are agnostic to the corruption in the sense that the algorithm does not depend on any knowledge of $\{c_{t}\}_{t=1}^{T}$.
4 The Intrinsic Limit of Learning from Imperfect Human Feedback
To understand the inherent difficulty of LIHF, in this section we develop regret lower bounds for a setting that is even easier to learn – that is, learning from a $\rho$-imperfect user (Definition 1) with a known $\rho$. While this non-agnostic may appear simpler, our main result of this section exhibits a tight lower bound of $\Omega(d\max\{\sqrt{T},T^{\rho}\})$ for strongly concave user utilities. We also show a similar lower bound hold for linear user utility as well. Our lower bound result differs in two key aspects from various previous lower bound results for online learning under corruptions (Agarwal et al., 2021; Bogunovic et al., 2020). Firstly, previous lower bound proofs require the adversaries’ capability to corrupt the reward feedback in an arbitrary way, whereas the corruption in our proof will follow Definition 1. It is unclear whether such strictly more restrictive corruption will be more tractable since it does offer the learner much more accurate estimation about the amount of corruption at each round. Secondly, previous construction of hard instances that attains regret lower bound relies on linear reward, whereas our lower bound holds for strongly concave utility functions, since they more realistically model human’s risk attitude and consumption utilities (Chiappori and Rochet, 1987). Strongly concave utility functions are generally perceived as easier to optimize and to learn, hence it is unclear whether it is also the case under imperfect user feedback. While both differences above seem to hint that our problem might be easier than previous online learning problem under corruption, we show that the answer is actually “No”.
Theorem 1 (A Tight Lower Bound).
There exists an LIHF instance with a $\rho$-imperfect user (Def. 1) and strongly concave user utility such that any learner has to suffer $\text{Reg}_{T}\geq\Omega(d\max\{\sqrt{T},T^{\rho}\})$ even with the knowledge of $\rho$. Moreover, this regret lower bound is tight, up to logarithmic terms. That is, for the same setting, there exists an online learning algorithm with $\text{Reg}_{T}\leq\tilde{O}(d\max\{\sqrt{T},T^{\rho}\})$.
This theorem reveals the fundamental limits of learning from $\rho$-imperfect users, even when the learner knows $\rho$. The core new lemma needed for proving the regret lower bound in Theorem 1 is to show that even in standard online learning with noisy bandit feedback about any pulled arm’s reward (as opposed to comparative feedback in dueling setups), the same regret lower bound holds. This is summarized in the following Lemma 1. We then convert this lower bound for standard bandit reward feedback to our setting of dueling/comparative feedback. Notably, Ye et al. (2024b) recently showed a matching upper bound for the standard bandit reward feedback setting assuming arbitrary corruption with total amount $C$. Hence Lemma 1 and their results together echoes our key insight in this section — that is, learning from $\rho$-imperfect user is as hard as learning from arbitrarily corrupted rewards in classic bandit reward feedback setting as well.
For linear user utility, perhaps unsurprisingly, we are able to show the same lower bound as follows (see Prop. 1). En route, we similarly show that the problem of learning directly from bandit linear reward feedback also suffers the same regret lower bound $\Omega(d\max\{\sqrt{T},T^{\rho}\})$. This happens to resolve an open problem proposed by Yao et al. (2022a) and implies that the regret upper bound shown in (Yao et al., 2022a) for the “learning from learning user” problem is tight in $T$.
Proposition 1.
There exists an LIHF instance with a $\rho$-imperfect user and linear user utility such that any learner knowing $\rho$ has to suffer $\text{Reg}_{T}\geq\Omega(d\max\{\sqrt{T},T^{\rho}\})$.
We defer the full proof of Theorem 1 and Proposition 1 to Appendix B.1 and B.2. Below we only state a representative core lemma for the strongly concave utility function case and its proof sketch, which is the crux of our lower bound proof. Notably, the proof for the matching upper bound in Theorem 1 is also highly non-trivial. Our starting point is the recent NC-SMD algorithm. However, to achieve the tight regret upper bound, we have to employ a novel and highly involved induction argument to identify the optimal learning rate. Through this process, we established a new connection between the convergence of the $L_{2}$-norm of the action $\|a_{t}-a^{*}\|_{2}$ and the regret itself (It has been experimentally verified at Appendix D.1.2).
Lemma 1.
Assume that the action space $\mathcal{A}$ is contained in a $d$-dimensional unit ball. Consider the utility function $\mu$ in the form $\mu_{\theta}(a)=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$, where $\theta\in\mathbb{R}^{d},\|\theta\|_{2}\leq 1$, is a random vector. Under corruption induced by $\rho$-imperfect user (Def. 1), for any fixed number of rounds $T$ and $d\leq\frac{1}{C_{\kappa}}T^{1-\rho}$, there exists a preference parameter $\theta$ such that for any learner $\mathcal{L}$ under reward feedback, even with the knowledge of $\rho$, has to suffer regret $\text{Reg}_{T}:=\mathbb{E}\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\}\geq\frac{d}{4}C_{\kappa}T^{\rho}.$
Proof (Proof Sketch of Lemma 1). We aim to show that there exists a corruption strategy such that the regret incurred by any learner $\mathcal{L}$ with corrupted reward feedback is no less than $\Omega(dT^{\rho})$. The formulation of such a corruption strategy hinges on the observation that the regret incurred by $\mathcal{L}$ is bounded below by the sum of the Kullback-Leibler (KL) divergence between the distribution of the corrupted reward feedback conditioned on different possible values of $\theta$. Specifically, let $\theta$ be uniformly drawn from $\{-\beta,\beta\}^{d}$. Given an action $a_{t}$, conditioned on $\theta_{i}>0$, the corrupted reward feedback $\hat{v}_{t}$ is
$\hat{v}_{t}=\mu_{\theta}(a_{t})+c_{t}(a_{t}|\theta_{i}>0)+\xi_{a_{t}}=\underbrace{\left(-\frac{1}{2}\|a_{t}\|^{2}+\sum_{j\neq i}\theta_{j}a_{t,j}\right)+\beta a_{t,i}+c_{t}(a_{t}|\theta_{i}>0)}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mu_{1}}}+\xi_{a_{t}}.$ (2)
Conditioned on $\theta_{i}<0$, the corrupted reward feedback is
$\hat{v}^{\prime}_{t}=\mu_{\theta}(a_{t})+c_{t}(a_{t}|\theta_{i}<0)+\xi_{a_{t}}=\underbrace{\left(-\frac{1}{2}\|a_{t}\|^{2}+\sum_{j\neq i}\theta_{j}a_{t,j}\right)-\beta a_{t,i}+c_{t}(a_{t}|\theta_{i}<0)}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mu_{2}}}+\xi_{a_{t}}.$ (3)
Consider the following corruption strategy which sets $c_{t}(a_{t}|\theta_{i}>0)=-\beta a_{t,i}$ and $c_{t}(a_{t})=\beta a_{t,i}$ to minimizes the KL divergence between $\hat{v}_{t}$ and $\hat{v}^{\prime}_{t}$. Together with $1$-strongly concavity of $\mu_{\theta}$, we can establish
$\mathbb{E}\left\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\right\}\geq\frac{1}{2}\max\left\{\sum^{T}_{t=1}(\frac{C_{\kappa}T^{\rho-1}}{\beta}-\beta)^{2},\frac{dT\beta^{2}}{2}\right\}.$ (4)
Since $d\leq\frac{1}{C_{\kappa}}T^{1-\rho}$, we can set $\beta=\sqrt{C_{\kappa}}T^{\frac{\rho-1}{2}}$ to satisfy $\|\theta\|_{2}\leq 1$ and to optimize the lower bound in Equation (4). Then we obtain $\mathbb{E}\left\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\right\}\geq dC_{\kappa}T^{\rho}/4$. Notice that the corruption level each round is bounded by $\beta\|a_{t}\|_{\infty}$. In the scenario when the radius of the action space $\mathcal{A}$ is upper bounded by $\beta$, the corruption budget each round is bounded by $C_{\kappa}t^{\rho-1}$. It implies that the corruption is induced by $\rho$-imperfect user, completing the proof.
5 The Efficiency-Robustness Tradeoff in LIHF
Section 4 shows that a lower bound regret of $\Omega(d\max\{\sqrt{T},T^{\rho}\})$ is intrinsic for learning from $\rho$-imperfect user, even with known $\rho$. In this section, we study what would be possible — but in a significantly more challenging learning setting with arbitrary and agnostic corruption. The reason for fighting for such generality is due to practical considerations; in reality, it is rarely possible to truly predict users’ irrationality or behaviors. Technique-wise, gradient-based methods are the most commonly used for addressing continuous action spaces, and so are our approaches. The key novelty in our results is not about the development of fundamentally new algorithmic techniques, but rather the conduct of more fine-grained analysis about two popular gradient-based methods — Dueling Bandit Gradient Descent (DBGD (Yue and Joachims, 2009)) and Noisy Comparison-based Stochastic Mirror Descent (NC-SMD) (Kumagai, 2017) — and the proof of how these algorithms can be “tailored” to achieve robustness against corruption. On a conceptual level, our key insight is that one can carefully tune the learning rates of these gradient-based methods to gracefully trade off learning efficiency for adversarial robustness — a interesting concept which we coin the efficiency-robustness tradeoff. Notably, en route developing our result, we also pin down a tight lower bound of $\Omega(T^{3/4})$ for the well-known DBGD algorithm which surprisingly remained an open question to date despite extensive studies of this algorithm.
Our finding reveals the intrinsic tradeoff between robustness and efficiency for gradient-based algorithms under agnostic corruption (Theorem 2). In general, the tradeoff is controlled by the exploration size in Algorithm 1 (depends on $\alpha$) (experimentally verified at Appendix D.1.4). A higher tolerance of agnostic corruption comes at a cost of worse regret in scenario without corruption, shown by Figure 1.
Input: Exploration rate $\delta$, exploitation rate $\gamma$, initial action $a_{1}=\mathbf{0}\in\mathcal{A}$.
1 for $t\in[T]$ do
2       Sample unit vector $u_{t}$ uniformly and set $a^{\prime}_{t}:=\mathcal{P}_{\mathcal{A}}\left(a_{t}+\delta u_{t}\right)$.
3       Present action pair $(a_{t},a^{\prime}_{t})$ to user and receive corrupted dueling feedback ^$\hat{\mathcal{F}}(a^{\prime}_{t},a_{t})$.
4       Compute gradient $\hat{g}_{t}:=-\frac{d}{\delta}\hat{\mathcal{F}}(a^{\prime}_{t},a_{t})u_{t}$.
5       Set step size $\eta:=\frac{\gamma\delta}{d}$ and update $a_{t+1}:=\mathcal{P}_{\mathcal{A}}(a_{t}-\eta\hat{g}_{t})$.
Algorithm 1 Dueling Bandit Gradient Descent (DBGD)
Theorem 2 (Efficiency-Robustness Tradeoff in DBGD).
For any $\alpha\in(0,\frac{1}{4}]$ and any number of round $T$ satisfying $T>\left(\frac{\sqrt{2Rd}L_{\mu}L_{2}}{\sqrt{13L}L_{\sigma}}\right)^{4}$, Algorithm 1 parameterized with $\gamma=\frac{R}{\sqrt{T}}$ and $\delta=\frac{\sqrt{2Rd}}{\sqrt{13L}T^{\alpha}}$ achieves $\text{Reg}_{T}\leq O(\sqrt{d}T^{1-\alpha}+\sqrt{d}T^{\alpha}C)$ for any total corruption level $C$.
At the core of this proof is a novel development of a regret decomposition lemma (Lemma 2), where we decompose the regret into two parts: regret of decision making under uncertainty and regret due to feedback imperfection. This decomposition has been proven to be useful for controlling regret of interactive decision-making in reinforcement learning Foster et al. (2023). However, to the best of our knowledge, Lemma 2 is the first to exhibit such a decomposition for dueling feedback, which contains much sparser information than the direct reward feedback as in RL or classic online learning in the aforementioned studies. Hence we believe the lemma may be of independent interest for future research on dueling bandits under corruption or erroneous observation. Our proof employs new techniques and is also more involved. Previous proofs of the decomposition lemma for reward feedback crucially hinges on an online estimation oracle Foster et al. (2023), which however cannot be constructed in our problem with dueling feedback with agnostic corruption. We thus have to develop new proof ideas tailored specifically for dueling bandit feedback. We start from quantifying the bias $b_{t}$ caused by adversarial corruption in gradient estimation, and then show that we can control the cumulative impact of bias by the exploration size $\delta$, i.e. $\sum^{T}_{t=1}\mathbb{E}\left(b_{t}^{\top}(a_{t}-a^{*})\right)\leq 2RdL_{\sigma}C/\delta$. Consequently, Lemma 2 shows that the regret of decision and observation error are jointly controlled by the exploration size $\delta$. We choose $\delta$ in order of $\Theta(T^{-\alpha}),\alpha\in(0,1/4]$ to achieve a balance between optimizing the learning objective (1) and controlling the gradient bias arising from the corrupted dueling feedback. Following this, we show $\text{Reg}_{T}$ is upper bounded by $O(\sqrt{d}T^{1-\alpha}+\sqrt{d}T^{\alpha}C)$.
Figure 1: Efficiency-Robustness Tradeoff
Lemma 2 (Regret Decomposition).
Let $b_{t}:=\frac{d}{\delta}\left(\hat{\mathbb{P}}(a^{\prime}_{t}\succ a_{t})-\mathbb{P}(a^{\prime}_{t}\succ a_{t})\right)u_{t},u_{t}\sim\mathbb{S}$, and choose $\lambda=\frac{L_{\sigma}}{L_{\sigma}-\delta L_{\mu}L_{\sigma}}$, we have
$\text{Reg}_{T}$ $\leq\underbrace{\lambda\left(\frac{R^{2}d}{\gamma\delta}+\frac{T\gamma d}{\delta}+13\delta LT\right)}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Regret\hskip 2.84526ptof\hskip 2.84526ptDecision}}+\underbrace{2\lambda\sum^{T}_{t=1}\mathbb{E}\left(b_{t}^{\top}(a_{t}-a^{*})\right)}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Observation\hskip 2.84526ptError}}.$
Note that Algorithm 1 works for general concave user utility. However, if the user displays a strongly concave utility, we can tune learning rate of NC-SMD to get better tradeoff. Specifically, we can get better regret upper bound $\tilde{O}(d\sqrt{T})$ in scenario without corruption at the same time affording $O(T^{3/4})$ agnostic corruption (see Prop 2). The proof of Proposition 2 shares similar structure as the proof Theorem 2, however working with a different algorithm. We direct readers to Appendix C.1 and C.4 to see full statement of Theorem 2, details of the Algorithm NC-SMD and proof of Proposition 2.
Proposition 2 (Efficiency-Robustness Tradeoff in NC-SMD).
For any $\alpha\in[\frac{1}{2},1)$, if the utility function is strongly concave, by choosing learning rate $\eta=\frac{\sqrt{\log T}}{2d}T^{-\alpha}$, for the Algorithm NC-SMD, we have $\text{Reg}_{T}\leq\tilde{O}(dT^{\alpha}+\sqrt{d}T^{\frac{1}{2}(1-\alpha)}C+dC)$ for any total corruption level $C$.
Figure 2: DBGD Lower Bound
An interesting corollary of Theorem 2 is that we can use it to show the DBGD upper bound proved in (Yue and Joachims, 2009) is tight, which to the best of our knowledge, has not been realized in the previous literature. In the following we provide a sketch of Corollary 1. It has been verified by experiments (see Figure 2, simulation details in Appendix D.1.1). We direct readers to Appendix C.2 for proof details.
Corollary 1 (Minimax Lower Bound for DBGD).
There exists a linear utility function $\mu$ and a link function $\sigma$ such that choosing $\alpha=1/4$, $\text{Reg}_{T}$ suffered by Algorithm 1 is at least $\Omega\left(T^{3/4}\right)$.
Proof  (Proof Sketch of Corollary 1). Corollary 1 shows that the standard DBGD algorithm with $\alpha=\frac{1}{4}$ can afford $O(T^{3/4})$ agnostic corruption. Moreover, we observe that we can make the sub-optimal arm and the optimal arm indistinguishable by using the total corruption level in the same order of the number of times sub-optimal arm that has been pulled. Therefore, there must exist an problem instance such that the sub-optimal arm has been pulled at least $\Theta(T^{3/4})$ times by DBGD. If not, we can create the parallel world to make DBGD suffer linear regret in one of the environments by using total corruption level less than $\Theta(T^{3/4})$, which contradicts Theorem 2. Intuitively, it says that if an environment is ”easy” to attack, the agnostic corruption level which an algorithm could afford can not exceed the worst possible regret suffered by the algorithm in scenario without corruption.
6 Experiments
In this section, we study the empirical performance of DBGD and NC-SMD to validate our theoretical analysis by running simulations on both synthetic datasets and real-world recommendation data in comparison with several baseline algorithms. Experimental details can be found at Appendix D.
6.1 Synthetic Data
Experiment Setup. We consider a standard experiment setup, which adopts a strongly concave utility $\mu_{\theta}(a):=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$, a logistic link function $\sigma(x)=\frac{1}{1+\exp(-x)}$. We choose $d=5$, and $T=10^{5}$. Our action space $\mathcal{A}$ is a $d$-dimensional ball with radius $R=10$. The preference parameter $\theta$ is randomly sampled from the surface of $\mathcal{A}$. In our problem setting, the optimal action $a^{*}$ is $\theta$ and $\mu_{\theta}(a^{*})=50$. We simulate agnostic arbitrary corruption for $C=T^{\rho},\rho\in[0.5,0.75]$. Specifically, we force the user to submit her least preferred item to the algorithm over the first $C$ rounds (for experiments on corruption induced by $\rho$-imperfect user, see Appendix D.1.3). For each value of $\rho$, we repeat the experiments for $5$ times, each under a different seed. We use Doubler and Sparring proposed by Ailon et al. (2014) as the baseline algorithms for comparison. We set bandit gradient descent (Flaxman et al., 2004) as the black-box bandit algorithm.
Figure 3: We show the log-log plot for regret occurred by different algorithms under arbitrary agnostic corruption. For each $\rho$, we present the line plot of average regret over $5$ simulations. The shaded region is the average regret $\pm$ one standard deviation. In addition, we fit the order of the average regret, i.e. $\mathbb{E}(\text{Reg}_{T})=T^{o}$, by least squares using last $1\%$ of data.
Results and Discussion. Figure 3 shows that standard DBGD ($\alpha=1/4$) and NC-SMD ($\alpha=1/2$) are able to tolerate $O(T^{3/4})$ unknown corruption, aligning with the theoretical result. Experimentally, NC-SMD enjoys the best performance under strongly concave utility. Empirically, Sparring has comparable performance as DBGD, even though it is a heuristic approach with no theoretical guarantee. The worst performance of Doubler might be attributed to the fact that its theoretical guarantee is only applicable to linear link function, which is different from our experiment setup.
6.2 Spotify Recommendation Data
Figure 4: Results on Spotify Recommendation
Evaluation Setup. We also evaluate our approach on Spotify recommendation data (Spotify, 2020). The objective is to recommend songs to incoming users. This dataset includes $17\times 10^{4}$ songs ($|\mathcal{A}|=17\times 10^{4}$), each is described by 15 distinct features (d = 15). We create 5 different user types by clustering the dataset into 5 groups. We use the average of the song embedding within each group as the user’s preference vector. Our aim is to recommend songs to users that have the highest cosine similarity (utility function) with their preference vectors. We rescale the cosine similarity to $[-100,100]$. We simulate corruption induced by $\rho$-imperfect user to model imperfect human feedback. We only test the performance of DBGD and Sparring because NC-SMD only works under strongly concave utility and Doubler only works under linear link function.
Results and Discussion. Figure 4 show that the performance of DBGD aligns with theoretical prediction on real-world recommendation data, even in the scenario when $\mathcal{A}$ is discrete and nonconvex. Both DBGD and Sparring are robust to agnostic corruption induced by $\rho$-imperfect user, however, empirically it seems that Sparring is invariant to $\rho$. This might because the song embeddings are highly sparse, which results in the fact that the proposed arms according to the heuristic approach are easy to distinguish, even under agnostic corruption.
Additional Experiments. We also test the performance of Versatile-DB (Saha and Gaillard, 2022), an approach robust to agnostic corruption with theoretical guarantee. However, this approach is not computationally efficient. When $K=2000$, running 100 iterations takes around 20 CPU hours and it’s performance is dominated by DBGD. We direct readers to Appendix D.2.1 to see the details.
7 Conclusion
In this paper, we model learning from imperfect human feedback as a corrupted continuous dueling bandit. We highlight the fundamental difficulty of LIHF by developing a regret lower bound. Then we show that by carefully tuning the learning rates of gradient-based methods, we can gracefully trade off learning efficiency for adversarial robustness. Our theoretical predictions also hold on real-world dataset. For future direction, it is interesting to generalize our setting to include contextual information and design algorithms robust to agnostic corruption. In addition, it is also an interesting open problem to design algorithms performs best-of-both-worlds, i.e. performing optimally both with and without agnostic corruption, in continuous action space.
A Connection between Imperfect User and Generalized Learnability
In this section, we establish the connection between adversarial corruption, corruption induced by generalized learnability, and corruption induced by $\rho$-imperfect user. Broadly speaking, holding total corruption budget $C$ constant, corruption induced by generalized learnability is a specific type of arbitrary corruption. Furthermore, if the proposed action pairs exhibit sufficient diversity, corruption induced by $\rho$-imperfect user is equivalent to corruption induced by generalized learnability.
In the following, we introduce the concept of the generalized learnability of the utility function $\mu$ in the scenario of pairwise comparison and the associated corruption induced throughout the user’s learning process.
Definition A.1 (Corruption induced by Generalized Learnability).
There exists an algorithm $\mathcal{G}$ such that, given any arbitrary $t$ pairs of actions $\{a_{s},a^{\prime}_{s}\}^{t}_{s=1}$ and the associated utility differences $\{y_{s}\}^{t}_{s=1}$, where $y_{s}:=\mu(a_{s})-\mu_{(}a^{\prime}_{s})$, $\mathcal{G}$ can output an estimated utility function $\mu_{t}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ of the true utility function $\mu:\mathbb{R}^{d}\rightarrow\mathbb{R}$ such that for any action pair $(a,a^{\prime})$ the following equation holds
$|c_{t}(a,a^{\prime})|=|\mu_{t}(a)-\mu(a)-(\mu_{t}(a^{\prime})-\mu(a^{\prime}))|\leq C_{0}\left(\|a-a^{\prime}\|^{2}_{\bar{V}_{t}^{-1}}\right)^{1-\rho}.$
$C_{0}$ is some positive constant, $\rho\in[0,1]$, and $\bar{V}_{t}:=\lambda\mathbb{I}_{d}+\sum^{t}_{s=1}(a_{s}-a^{\prime}_{s})(a_{s}-a^{\prime}_{s})^{\top}$, where $\lambda$ is some positive constant to ensure that $\bar{V}_{t}$ is full rank.
Def. A.1 is a variant of the generalized learnability assumption proposed by Wang et al. (2023) and Yao et al. (2022a). Essentially, the parameter $\rho$ can be viewed as the user’s irrationality level. A higher value of $\rho$ implies lower learning speed of the user’s utility function $\mu$, hence a larger magnitude of the total induced corruption. Let’s consider arbitrary corruption with total corruption budget $C=\Theta(T^{\rho})$. Using Lemma A.1, we can show that $c_{t}(a,a^{\prime})$ induced by the generalized learnability of $\mu$ satisfies $\sum^{T}_{t=1}|c_{t}(a_{t},a^{\prime}_{t})|\leq O(T^{\rho})$, implies that it is a specific type of arbitrary corruption.
In the scenario when $\lambda_{\min}(\bar{V}_{t})$ increases in the order of $\Theta(t)$, $(\|a-a^{\prime}\|^{2}_{\bar{V}^{-1}_{t}})^{1-\rho}$ has the order $\Theta(t^{\rho-1})$, which implies there exists a positive constant $C_{\kappa}$ that at round $t$, given arbitrarily proposed action pairs $(a,a^{\prime})$, the magnitude of corruption $|c_{t}(a,a^{\prime})|\leq C_{\kappa}t^{\rho-1}$, which means that corruption induced by a $\rho$-imperfect user coincides with corruption induced by the generalized learnability of $\mu$.
We would like to highlight that the condition which $\lambda_{\min}(\bar{V}_{t})$ increases in the order of $\Theta(t)$ could be satisfied by gradient-descent based algorithm. Take Algorithm 1 as an example. For simplicity, let’s assume $a^{*}$ is far away from the boundary of $\mathcal{A}$ and projection never happens. It implies that $a^{\prime}_{t}=a_{t}+\delta u_{t},\forall t\in[T]$. This assumption is without loss of generality, since most of the proposed action $a^{\prime}_{t}$ will belong to the interior of the action space $\mathcal{A}$. This is because we start with $a_{0}=0$ and has small exploration size $\delta\sim\Theta(T^{-\alpha})$. Roughly speaking, we can express $\bar{V}_{t}=\lambda\mathbb{I}_{d}+\delta^{2}\sum^{t}_{s=1}u_{s}u^{\top}_{s}$, where $u_{s}$ is uniformly sampled from $\mathbb{S}^{d}$. Then applying Lemma A.2, we have $\lambda_{\min}(\bar{V}_{t})\sim\Theta(t)$ with high probability.
Lemma A.1 (Generalized Elliptical Lemma in Wang et al. (2023)).
Suppose $V_{0}\in\mathbb{R}^{d\times d}$ is any positive definite matrix, $a_{1},\ldots,a_{T}\in\mathbb{R}^{d}$ is a sequence of vectors with bounded $l_{2}$ norm and $\bar{V}_{t}:=V_{0}+\sum^{t}_{s=1}a_{s}a^{\top}_{s}$. Then for any $\rho\in[0,1]$, the following inequality holds with probability at least $1-\delta$
$\sum^{T}_{t=1}\left(\min\left\{1,\|a_{t}\|^{2}_{\bar{V}^{-1}_{t}}\right\}\right)^{1-\rho}\leq 2^{1-\rho}T^{\rho}\log^{1-\rho}\left(\frac{\det(\bar{V}_{t})}{\det(V_{0})}\right).$
Lemma A.2 (Proposition 1 in Li et al. (2017)).
Define $V_{t}:=\sum^{t}_{s=1}u_{s}u^{\top}_{s}$, where $u_{s}$ is drawn iid from some distribution $\nu$ with support on the unit ball, $\mathbb{B}^{d}$. Furthermore, let $\Sigma:=\mathbb{E}[u_{s}u^{\top}_{s}]$ be the second moment matrix. $B$ and $\delta>0$ are two positive constants. Then, there exist positive, universal constants $C_{1}$ and $C_{2}$ such that $\lambda_{\min}(V_{s})\geq B$ with probability at least $1-\delta$, as long as
$t\geq\left(\frac{C_{1}\sqrt{d}+C_{2}\sqrt{\log(1/\delta)}}{\lambda_{\min}(\Sigma)}\right)^{2}+\frac{2B}{\lambda_{\min}(\Sigma)}.$
B Missing Proofs in Theorem 1 and Proposition 1
B.1 Proof for Theorem 1: A Tight Lower Bound
Theorem (Theorem 1 restated).
There exists an LIHF instance with a $\rho$-imperfect user (Def. 1) and strongly concave user utility such that any learner has to suffer $\text{Reg}_{T}\geq\Omega(d\max\{\sqrt{T},T^{\rho}\})$ even with the knowledge of $\rho$. Moreover, this regret lower bound is tight, up to logarithmic terms. That is, for the same setting, there exists an online learning algorithm with $\text{Reg}_{T}\leq\tilde{O}(d\max\{\sqrt{T},T^{\rho}\})$.
Proof (Proof for Regret Lower Bound in Theorem 1). Before showing the regret lower bound proof for Theorem 1, let’s first establish the following Lemma B.1, which builds the connection between regret occurred under bandit reward feedback and dueling feedback. This proof is inspired by Theorem 6 in Yao et al. (2022a). To prove Lemma B.1, in addition to $\text{Reg}_{T}$, we introduce a new metric to measure the performance of algorithm under dueling feedback, which we coin functional regret $\text{Reg}^{\text{FO}}_{T}$, defined as follows.
$\text{Reg}^{\text{FO}}_{T}:=\mathbb{E}\left\{\sum^{T}_{t=1}\mu(a^{*})-\mu(a_{t})+\mu(a^{*})-\mu(a^{\prime}_{t})\right\}.$ (5)
Lemma B.1.
Given a utility function $\mu$, if the regret lower bound for algorithm with bandit reward feedback is $\overline{\text{Reg}}$, then any learner $\mathcal{L}$ with dueling feedback has to occur regret $\text{Reg}^{\text{FO}}_{T}\geq 2\overline{\text{Reg}}$.
Proof  (Proof for Lemma B.1). We prove our claim by contradiction. Let $(a_{0,t},a_{1,t})$ be the pair of recommendation at round $t$. Suppose $\text{Reg}^{\text{FO}}_{T}<2\overline{\text{Reg}}$. As a result, at least one of the following inequality must hold:
$\begin{split}\mathbb{E}\left\{\sum^{T}_{t=1}f(a_{0,t})-f(a^{*})\right\}&\leq\overline{\text{Reg}};\\ \mathbb{E}\left\{\sum^{T}_{t=1}f(a_{1,t})-f(a^{*})\right\}&\leq\overline{\text{Reg}}.\end{split}$ (6)
Consider a principal who can observe the interaction between a user and the learner $\mathcal{L}$, then Then we can construct two algorithms $\mathcal{L}_{0}$ and $\mathcal{L}_{1}$ as follows.
Input: the time horizon $T$
1 for $t\leq T$ do
2       Call algorithm $\mathcal{L}$ to generate two candidates $(a_{0,t},a_{1,t})$.
3       Present $(a_{0,t},a_{1,t})$ to user and received the feedback.
4       Return the feedback $a_{*,t}$ to algorithm $\mathcal{L}$ and update $\mathcal{L}$.
Output : the sequential decision $\{a_{i,t}\}^{T}_{t=1}$
Algorithm 2 Algorithm $\mathcal{L}_{i}$
From Equation (6), we know that at least one of $\{\mathcal{L}_{0},\mathcal{L}_{1}\}$ achieves an expected regret lower than $\overline{\text{Reg}}$. However, we know all algorithm with bandit feedback with utility function $\mu$ has to occur regret at least $\overline{\text{Reg}}$, contradiction. Therefore, Lemma B.1 must hold, which completes the proof.
At the next step, we will prove Theorem 1 by discussing the value of $\rho$ case by case. Consider the scenario when $\rho\leq 0.5$. This is equivalent to say the total corruption level $C\leq\sqrt{T}$. In this scenario, since the utility function is strongly concave, applying Lemma B.2, we know $\overline{\text{Reg}}=0.02\min\left\{T,d\sqrt{T}\right\}$. By Lemma B.1, we get $\text{Reg}^{\text{FO}}_{T}\geq 0.04\min\left\{T,d\sqrt{T}\right\}$. Consider a linear link function $\sigma(x)=\frac{1}{2}+\frac{1}{2}x$, we have $\text{Reg}_{T}\geq 0.02\min\left\{T,d\sqrt{T}\right\}\geq\Omega(d\max\{\sqrt{T},T^{\rho}\})$, which completes the proof.
Lemma B.2 (Theorem 6 in (Shamir, 2013)).
Let the number of rounds $T$ be fixed. The for any learner $\mathcal{L}$, there exists a quadratic function of the form $\mu_{\theta}(a)=\theta^{\top}a-\frac{1}{2}\|a\|^{2}$ which is minimized at $\theta$ and $\|\theta\|_{2}\leq 0.5$ such that
$\mathbb{E}\left(\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\right)\geq 0.02\min\left\{T,d\sqrt{T}\right\}.$
Now, let’s focus on the scenario when $\rho>0.5$, which is equivalent to $C>\sqrt{T}$. In essence, we want to extend Lemma B.2 to the scenario in presence of adversarial corruption induced by $\rho$-imperfect user. If we can show the regret lower bound can be generalized to $\Omega(dT^{\rho})$, then applying the same technique which uses Lemma B.1 to connect regret suffered under bandit reward feedback and regret suffered under dueling feedback and choose linear link function to connect $\text{Reg}^{\text{FO}}_{T}$ and $\text{Reg}_{T}$, we will get the desired regret lower bound. The extension of Lemma B.2 is proven at section B.1.1.
B.1.1 Proof for Lemma 1: Regret Lower Bound for Strongly Concave Utility
Lemma (Lemma 1 restated).
Assume that the action space $\mathcal{A}$ is contained in a $d$-dimensional unit ball. Consider the utility function $\mu$ in the form $\mu_{\theta}(a)=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$, where $\theta\in\mathbb{R}^{d},\|\theta\|_{2}\leq 1$ is a random vector. Under corruption induced by $\rho$-imperfect user (Def. 1), for any fixed number of rounds $T$ and $d\leq\frac{1}{C_{\kappa}}T^{1-\rho}$, there exists a preference parameter $\theta$ such that for any learner $\mathcal{L}$ under reward feedback, even with the knowledge of $\rho$, has to suffer regret $\text{Reg}_{T}:=\mathbb{E}\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\}\geq\frac{d}{4}C_{\kappa}T^{\rho}.$
Proof  (Proof for Lemma 1). To prove Lemma 1, we aim to prove that there exists a corruption strategy such that the regret incurred by any learner $\mathcal{L}$ with corrupted dueling feedback is not less than $\Omega(dT^{\rho})$. The formulation of such a corruption strategy hinges on the observation that when the utility function $\mu_{\theta}$ adopts the quadratic form parameterized by $\theta$, specifically $\mu_{\theta}(a)=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$, which is both smooth and strongly concave, then the regret incurred by $\mathcal{L}$ is bounded below by the sum of the Kullback-Leibler (KL) divergence between the distribution of the corrupted reward feedback obtained at round $t$, denoted $\hat{v}_{t},t\in[1,T]$, conditioned on different possible values of $\theta$. (see Lemma B.4). Assume that $\theta$ is uniformly drawn from $\{-\beta,\beta\}^{d}$, given an action $a_{t}$, conditioned on $\theta_{i}>0$, the corrupted reward feedback $\hat{v}_{t}$ is
$\hat{v}_{t}=\mu_{\theta}(a_{t})+c_{t}(a_{t}|\theta_{i}>0)+\xi_{a_{t}}=\underbrace{\left(-\frac{1}{2}\|a_{t}\|^{2}+\sum_{j\neq i}\theta_{j}a_{t,j}\right)+\beta a_{t,i}+c_{t}(a_{t}|\theta_{i}>0)}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mu_{1}}}+\xi_{a_{t}}.$
Conditioned on $\theta_{i}<0$, the corrupted reward feedback $\hat{v}_{t}$ is
$\hat{v}_{t}=\mu_{\theta}(a_{t})+c_{t}(a_{t}|\theta_{i}<0)+\xi_{a_{t}}=\underbrace{\left(-\frac{1}{2}\|a_{t}\|^{2}+\sum_{j\neq i}\theta_{j}a_{t,j}\right)-\beta a_{t,i}+c_{t}(a_{t}|\theta_{i}<0)}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mu_{2}}}+\xi_{a_{t}}.$
We use $c_{t}(a_{t}|\theta_{i}>0)$ and $c_{t}(a_{t}|\theta_{i}<0)$ to empathize the fact that the magnitude and sign of corruption can be dependent on $\theta$. $\xi_{a_{t}}$ follows a standard Gaussian distribution. Therefore, $\hat{v}_{t}$ in Equation 2 follows $\text{N}(\mu_{1},1)$. $\hat{v}_{t}$ in Equation 3 follows $\text{N}(\mu_{2},1)$. By using Lemma B.3, we have
$\mathbb{D}_{t,i}=\text{D}_{\text{KL}}(\text{N}(\mu_{1},1)||\text{N}(\mu_{2},1))=(\mu_{1}-\mu_{2})^{2}.$
Lemma B.3 (KL Divergence for Normal Distribution).
Let $N(\mu,\sigma^{2})$ represent a Gaussian distribution variable with mean $\mu$ and variance $\sigma^{2}$. Then
$\text{D}_{\text{KL}}(N(\mu_{1},\sigma^{2})||N(\mu_{2},\sigma^{2}))=\frac{(\mu_{1}-\mu_{2})^{2}}{2\sigma^{2}}.$
To optimize the lower bound in Lemma B.4, the adversary selects $c_{t}(a_{t}|\theta_{i}>0)$ and $c_{t}(a_{t}|\theta_{i}<0)$ to minimize the difference between $\mu_{1}$ and $\mu_{2}$. Consider the following corruption strategy: when $\theta_{i}>0$, set $c_{t}(a_{t}|\theta_{i}>0)=-\beta a_{t,i}$; when $\theta_{i}<0$, set $c_{t}(a_{t}|\theta_{i}<0)=\beta a_{t,i}$. This strategy ensures that $\mu_{1}=\mu_{2}$.
Next step is to determine the value of $\beta$. Notice that the corruption budget is $C_{\kappa}T^{\rho}$. To execute the corruption strategy described above, it requires
$\sum^{T}_{t=1}|c_{t}(a_{t})|\leq\beta\sum^{T}_{t=1}\|a_{t}\|_{\infty}\leq C_{\kappa}T^{\rho}.$ (7)
Therefore, by exploiting the fact that $\mu_{\theta}$ is $1$-strongly concave, we establish another lower bound for the regret by
$\mathbb{E}\left\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\right\}$ $\geq\frac{1}{2}\mathbb{E}\left\{\sum^{T}_{t=1}\|a_{t}-\theta\|^{2}_{2}\right\}$
$\geq\frac{1}{2}\mathbb{E}\left\{\sum^{T}_{t=1}(\|a_{t}\|_{\infty}-\beta)^{2}\right\}$
$\geq\frac{1}{2}\sum^{T}_{t=1}(C_{\kappa}T^{\rho-1}/\beta-\beta)^{2}.$
Since $|a_{t,i}-\theta_{i}|\geq||a_{t,i}|-\beta|$ for all $i$, we obtain the second inequality. By utilizing the corruption constraint (7), we derive the last inequality, which is minimized when $\|a_{t}\|_{\infty}=\frac{C_{\kappa}T^{\rho-1}}{\beta}$ for all $t$. Together with Lemma B.4, we have:
$\mathbb{E}\left\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\right\}\geq\frac{1}{2}\max\left\{\sum^{T}_{t=1}(C_{\kappa}T^{\rho-1}/\beta-\beta)^{2},\frac{dT\beta^{2}}{2}\right\}.$ (8)
We select $\beta$ to optimize the lower bound in Equation (8). Since $d\leq\frac{1}{C_{\kappa}}T^{1-\rho}$, we can set $\beta=\sqrt{C_{\kappa}}T^{\frac{\rho}{2}-\frac{1}{2}}$ to satisfy $\|\theta\|_{2}\leq 1$. Then we obtain $\mathbb{E}\left\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\right\}\geq dC_{\kappa}T^{\rho}/4$. Notice that the corruption level each round is bounded by $\beta\|a_{t}\|_{\infty}$. Consider the scenario when the radius of the action space $\mathcal{A}$ is upper bounded by $\beta$, then the corruption budget each round is bounded by $C_{\kappa}t^{-1+\rho}$, which satisfies the corruption induced by $\rho$-imperfect user, which completes the proof.
Lemma B.4.
Let’s consider the utility function $\mu_{\theta}(a)=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$. Let $\hat{v}_{1},\hat{v}_{2},\ldots,\hat{v}_{T}$ be a sequence of corrupted reward feedback obtained by a learner $\mathcal{L}$. Then there exists a $\theta\in\mathbb{R}^{d},\|\theta\|_{2}\leq 1$, uniformly drawn from $\{-\beta,\beta\}^{d}$, such that the regret occurred by $\mathcal{L}$ is
$\mathbb{E}\left\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\right\}\geq\frac{dT\beta^{2}}{4}\left(1-\sqrt{\frac{1}{d}\sum^{d}_{i=1}\sum^{T}_{t=1}\mathbb{D}_{t,i}}\right).$
$\mathbb{D}_{t,i}:=\sup_{\{\theta_{j}\}_{j\neq i}}\text{D}_{\text{KL}}\left(\mathbb{P}(\hat{v}_{t}|\theta_{i}>0,\{\theta_{j}\}_{j\neq i},\{\hat{v}_{l}\}^{t-l}_{l=1})||\mathbb{P}(\hat{v}_{t}|\theta_{i}<0,\{\theta_{j}\}_{j\neq i},\{\hat{v}_{l}\}^{t-l}_{l=1})\right)$. $\text{D}_{\text{KL}}$ is the KL divergence between two distributions.
Proof  (Proof for Lemma B.4). Using the similar argument in Shamir (2013), assume that the leaner $\mathcal{L}$ is deterministic: $a_{t}$ is a deterministic function of the realized corrupted reward feedback $\hat{v}_{1},\hat{v}_{2},\ldots,\hat{v}_{t-1}$ at $a_{1},a_{2},\ldots,a_{t-1}$. This assumption is without loss of generality, since any random learners can be seen as a randomization over deterministic learning algorithms. Thus a lower bound which holds uniformly for any deterministic $\mathcal{L}$ would also hold over a randomization. To lower bound Equation (9), we use Lemma B.5, which relates this to the question of how informative are the query values (as measured by Kullback-Leibler divergence) for determining the sign of $\theta$’s coordinates. Intuitively, the more similar the query values are, the smaller is the KL divergence and the harder it is to distinguish the true sign of $\theta_{i}$, leading to a larger lower bound. In addition, we are facing a powerful adversary who has the complete knowledge of the problem and is able to add corruption on the query value to make they are even more similar, which resulting a even smaller KL divergence, consequently, an even larger lower bound. Let $\bar{a}_{T}=\frac{1}{T}\sum^{T}_{t=1}a_{t}$ represent the average action, we have
$\mathbb{E}\left\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\right\}$ $=T\mathbb{E}\left\{\frac{1}{T}\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\right\}$
$\geq T\mathbb{E}\left(\mu_{\theta}(a^{*})-\mu_{\theta}(\bar{a}_{T})\right)$
$\geq T\mathbb{E}\left(\frac{1}{2}\|\bar{a}_{T}-\theta\|^{2}\right)$
$=T\mathbb{E}\left(\frac{1}{2}\sum^{d}_{i=1}(\bar{a}_{i}-\theta_{i})^{2}\right)$
$\geq\mathbb{E}\left(\frac{\beta^{2}T}{2}\sum^{d}_{i=1}\mathbb{I}_{\bar{a}_{i}\theta_{i}<0}\right)$ (9)
$\geq\frac{dT\beta^{2}}{4}\left(1-\sqrt{\frac{1}{d}\sum^{d}_{i=1}\sum^{T}_{t=1}\mathbb{D}_{t,i}}\right).$
We get the second inequality by using the fact that $\mu_{\theta}$ is 1-strongly concave. We get the last inequality by using Lemma B.5, which completes the proof.
Lemma B.5 (Lemma 4 in (Shamir, 2013)).
Let $\theta$ be a random vector, none of those coordinates is supported on 00. Let $\hat{v}_{1},\hat{v}_{2},\ldots,\hat{v}_{T}$ be a sequence of values obtained by a deterministic learner $\mathcal{L}$ returning a point $\bar{a}_{T}$ (so that the action $a_{t}$ is a deterministic function of $\hat{v}_{1},\ldots,\hat{v}_{t-1}$ and $\bar{a}_{T}$ is a deterministic function of $\hat{v}_{1},\ldots,\hat{v}_{T}$). Then we have
$\mathbb{E}\left(\sum^{d}_{i=1}\mathbb{I}_{\bar{a}_{i}\theta_{i}}\right)\geq\frac{d}{2}\left(1-\sqrt{\frac{1}{d}\sum^{d}_{i=1}\sum^{T}_{t=1}\mathbb{D}_{t,i}}\right),$
where $\mathbb{D}_{t,i}=\sup_{\{\theta_{j}\}_{j\neq i}}\text{D}_{\text{KL}}\left(\mathbb{P}(\hat{v}_{t}|\theta_{i}>0,\{\theta_{j}\}_{j\neq i},\{\hat{v}_{l}\}^{t-l}_{l=1})||\mathbb{P}(\hat{v}_{t}|\theta_{i}<0,\{\theta_{j}\}_{j\neq i},\{\hat{v}_{l}\}^{t-l}_{l=1})\right)$, $\text{D}_{\text{KL}}$ represents the KL divergence between two distributions.
We defer reader to Appendix C.3 for the proof of matching regret upper bound and the detailed description of the algorithm which achieves that upper bound.
B.2 Proof for Proposition 1: Regret Lower Bound for Linear Utility
Proposition (Proposition 1 restated).
There exists an LIHF instance with a $\rho$-imperfect user and linear user utility such that any learner knowing $\rho$ has to suffer $\text{Reg}_{T}\geq\Omega(d\max\{\sqrt{T},T^{\rho}\})$.
Proof  (Proof for Proposition 1). The proof structure is very similar to the lower bound proof in Theorem 1. We start by discussing the value of $\rho$. Consider the scenario when $\rho\leq 0.5$. Applying Lemma B.6 and B.1 together with linear link function, we have $\text{Reg}_{T}\geq\Omega(d\max\{\sqrt{T},T^{\rho}\})$, which completes the proof.
Lemma B.6 ((Lattimore and Szepesvári, 2020)).
Let $\mathcal{A}=[-1,1]^{d}$ and $\Theta=[-T^{-\frac{1}{2}},T^{-\frac{1}{2}}]^{d}$. Consider the linear reward function $r_{t}=\theta^{\top}A_{t}+\epsilon_{t}$, $\epsilon_{t}$ is independent Gaussian noise with mean 0 variance 1. Then for any learner $\mathcal{L}$, there exists a vector $\theta\in\Theta$ such that
$\text{Reg}_{T}(\mathcal{A},\theta)\geq\frac{\exp(-2)}{8}d\sqrt{T}.$
Now, let’s focus on the scenario when $\rho>0.5$ and the essence is to extend Lemma B.6 to the scenario in presence of adversarial corruption induced by $\rho$-imperfect user, which is shown in Lemma B.7
Lemma B.7.
Assume that the action space $\mathcal{A}\subset[-1,1]^{d}$. Given corruption induced by $\rho$-imperfect user (Def. 1), for stochastic linear bandit, there exists a $\theta\in[-C_{\kappa}T^{\rho-1},C_{\kappa}T^{\rho-1}]^{d}$ such that any learner $\mathcal{L}$ even with the knowledge of $\rho$ has to suffer regret no less than $\frac{d}{8}C_{\kappa}T^{\rho}.$
Proof  (Proof for Lemma B.7). The proof extends Lemma B.6 to a scenario in presence of corruption. We want to construct a parameter family and a corruption strategy such that for all algorithm $\mathcal{L}$, it will occur at least $\Omega(dT^{\rho})$ regret. Consider the action set $\mathcal{A}\in[-1,1]^{d}$ and $\Theta=\{-\beta,\beta\}^{d}$. For any learner $\mathcal{L}$, we can lower bound its regret by
$\text{Reg}_{T}(\mathcal{A},\theta)$ $=\mathbb{E}_{\theta}\left[\sum^{T}_{t=1}\sum^{d}_{i=1}(\text{sign}(\theta_{i})-a_{ti})\theta_{i}\right]$
$\geq\beta\sum^{d}_{i=1}\mathbb{E}_{\theta}\left[\sum^{T}_{t=1}\mathbb{I}\{\text{sign}(a_{ti})\neq\text{sign}(\theta_{i})\}\right]$
$\geq\frac{T\beta}{2}\sum^{d}_{i=1}\mathbb{P}_{\theta}\left(\sum^{T}_{t=1}\mathbb{I}\{\text{sign}(a_{ti})\neq\text{sign}(\theta_{i})\}\geq\frac{T}{2}\right)$
Let’s denote
$p_{\theta_{i}}=\mathbb{P}_{\theta}\left(\sum^{T}_{t=1}\mathbb{I}\{\text{sign}(a_{ti})\neq\text{sign}(\theta_{i})\}\geq\frac{T}{2}\right)$
Let $i\in[d]$ and $\theta\in\Theta$ be fixed, and let $\theta^{\prime}_{j}=\theta_{j}$, for $j\neq i$ and $\theta^{\prime}_{i}=-\theta_{i}$. Then using Lemma B.8, we have
$p_{\theta_{i}}+p_{\theta^{\prime}_{i}}\geq\frac{1}{2}\exp(-\mathbb{D}(\mathbb{P}_{\theta},\mathbb{P}_{\theta^{\prime}}))=\frac{1}{2}\exp\left(-\mathbb{E}\left[\sum^{T}_{t=1}\text{D}_{\text{KL}}(P_{a_{t}},P^{\prime}_{a_{t}})\right]\right).$
$P_{a_{t}}$ is the distribution of corrupted reward observed by $\mathcal{L}$ after playing action $a_{t}$ when reward parameter is $\theta$. Similarly, $P^{\prime}_{a_{t}}$ is the distribution of corrupted observed by $\mathcal{L}$ after playing action $a_{t}$ when the reward parameter is $\theta^{\prime}$.
Lemma B.8 (Bretagnolle-Huber inequality).
Let $\mathbb{P}$ and $\mathbb{Q}$ be probability measures on the same measurable space $(\Omega,\mathcal{F})$. Let $A\in\mathcal{F}$ be any arbitrary event and $A^{c}$ is the complement of $A$. Then we have
$\mathbb{P}(A)+\mathbb{Q}(A^{c})\geq\frac{1}{2}\exp\left(-D_{\text{KL}}(\mathbb{P},\mathbb{Q})\right).$ (10)
In presence of adversarial corruption, for $\theta$, the corrupted reward is
$\hat{r}_{t}=\sum_{j\neq i}a_{tj}\theta_{j}+a_{ti}\theta_{i}+c_{t}(a_{t}|\theta)+\epsilon_{t}.$ (11)
For $\theta^{\prime}$, we have
$\hat{r}_{t}=\sum_{j\neq i}a_{tj}\theta_{j}-a_{ti}\theta_{i}+c_{t}(a_{t}|\theta^{\prime})+\epsilon_{t}.$ (12)
Consider the corruption strategy such that $c_{t}(a_{t}|\theta)=-a_{t,i}\theta_{i}$ in (11) and $c_{t}(a_{t}|\theta^{\prime})=a_{t,i}\theta_{i}$ in (12). This is achievable since we assume the adversary has complete knowledge of the problem instance. By doing so, the corrupted reward $\hat{r}_{t}$ are from the same distribution regardless whether the preference parameter is $\theta$ or $\theta^{\prime}$. This is because $\theta$ and $\theta^{\prime}$ only differs in the $i$-th coordinate with the magnitude $\beta$, and this difference could be masked by the corruption, resulting in $\mathbb{E}\left[\sum^{T}_{t=1}\mathbb{D}(P_{A_{t}},P^{\prime}_{A_{t}})\right]=0$, which makes it indistinguishable by $\mathcal{L}$. Notice that executing such a corruption strategy requires total corruption budget
$\sum^{T}_{t=1}|c_{t}(a_{t})|\leq\beta\sum^{T}_{t=1}\|a_{t}\|_{\infty}\leq C_{\kappa}T^{\rho}.$ (13)
Choose $\beta=C_{\kappa}T^{\rho-1}$, Equation (13) holds. Therefore, we have
$p_{\theta_{i}}+p_{\theta^{\prime}_{i}}\geq\frac{1}{2}.$
Applying an ”averaging hammer” over all $\theta\in\Theta$, which satisfies $|\Theta|=2^{d}$, we get
$\sum_{\theta\in\Theta}\frac{1}{|\Theta|}\sum^{d}_{i=1}p_{\theta_{i}}=\frac{1}{|\Theta|}\sum^{d}_{i=1}\sum_{\theta\in\Theta}p_{\theta_{i}}\geq\frac{d}{4}.$
This implies that there exists a $\theta\in\Theta$ such that $\sum^{d}_{i=1}p_{\theta_{i}}\geq\frac{d}{4}$. Therefore we have
$\text{Reg}_{T}(\mathcal{A},\theta)\geq\frac{dC_{\kappa}}{8}T^{\rho},$
which completes the proof. We want to highlight that the corruption level each round $|c_{t}(a_{t})|$ is bounded by $\beta\|a_{t}\|_{\infty}\leq C_{\kappa}T^{\rho-1}\leq C_{\kappa}t^{\rho-1},\forall t$, satisfying the corruption induced by $\rho$-imperfect user.
C Missing Proofs in Theorem 2 and Proposition 2
C.1 Proof for Theorem 2: Efficiency-Robustness Tradeoff in DBGD
In this section, we present the regret upper for dueling bandit gradient descent Yue and Joachims (2009) in presence of arbitrary and agnostic corruption. Before presenting the proof, we make several remarks.
First, in Yue and Joachims (2009), the utility function $\mu$ is assumed to be strictly concave to ensure the uniqueness of the maximizer. However, in scenarios where the maximizer is unique within the action set for a concave utility function $\mu$, the requirement for ”strictness” can be relaxed.
Second, we highlight the notation differences. We use $a$ to denote actions while Yue and Joachims (2009) uses $w$. In our proof, we define $P_{t}(a):=\sigma\left(\mu(a_{t})-\mu(a)\right)$ to represent the probability of the event $a_{t}\succ a$. While in Yue and Joachims (2009), it is denoted as $c_{t}(w)$. Just to highlight, its corrupted version $\hat{P}_{t}(a):=\sigma\left(\mu(a_{t})-\mu(a)+c_{t}(a_{t},a)\right)$ is newly introduced in our paper. Moreover, $\bar{P}_{t}(a):=\mathbb{E}_{x\in\mathbb{B}}\left[P_{t}(\mathcal{P}_{\mathcal{A}}(a+\delta x))\right]$ is denoted as $\hat{\epsilon}_{t}(w)+\frac{1}{2}$ in Yue and Joachims (2009). The definition of $\mathcal{F}\left(\mathcal{P}_{\mathcal{A}}\left(a_{t}+\delta u_{t}\right),a_{t}\right)$ is same as $X_{t}\left(\mathcal{P}_{W}(w_{t}+\delta u_{t})\right)$ defined in Yue and Joachims (2009), except different notation.
Theorem (Theorem 2 restated).
For any $\alpha\in(0,\frac{1}{4}]$ and any number of round $T$ satisfying $T>\left(\frac{\sqrt{2Rd}L_{\mu}L_{2}}{\sqrt{13L}L_{\sigma}}\right)^{4}$, Algorithm 1 parameterized with $\gamma=\frac{R}{\sqrt{T}}$ and $\delta=\frac{\sqrt{2Rd}}{\sqrt{13L}T^{\alpha}}$ achieves $\text{Reg}_{T}\leq O(\sqrt{d}T^{1-\alpha}+\sqrt{d}T^{\alpha}C)$ for any total corruption level $C$.
Proof  (Proof for Theorem 2). The proof of Theorem 2 builds on the proof of Yue and Joachims (2009), where we extend it to a setting in presence of agnostic corruption. The cornerstone of the theoretical analysis relies on developing the regret decomposition lemma (Lemma 2), where we novelly decompose the regret into two parts: regret of decision and observation error. This decomposition is achieved through quantifying the bias, $b_{t}$, caused by adversarial corruption, in gradient estimation (Lemma C.1). We further use Lemma C.5 to upper bound the observation error in terms of the exploration size $\delta$. Lemma 2 and Lemma C.5 reveal that the regret of decision and observation error are jointly controlled by the exploration size $\delta$. We choose $\delta$ in order of $O(T^{-\alpha})$ to achieve a balance between optimizing the learning objective (1) and controlling the gradient bias arising from the corrupted dueling feedback. Consequently, we have $\text{Reg}_{T}$ is upper bounded by $O(\sqrt{d}T^{1-\alpha}+\sqrt{d}T^{\alpha}C)$.
The development of the proof is based on the observation that we decompose bias in the gradient caused by corruption. Notice that without corruption, DBGD performs expected gradient ascent, where the gradient $g_{t}$ is defined as follows
$g_{t}=-\frac{d}{\delta}\mathcal{F}\left(\mathcal{P}_{\mathcal{A}}\left(a_{t}+\delta u_{t}\right),a_{t}\right)u_{t}.$
In the presence of adversarial corruption, the corrupted gradient is
$\hat{g}_{t}=-\frac{d}{\delta}\hat{\mathcal{F}}\left(\mathcal{P}_{\mathcal{A}}\left(a_{t}+\delta u_{t}\right),a_{t}\right)u_{t}.$
Therefore, we quantify the bias in the following lemma.
C.1.1 Proof for Lemma C.1: Gradient Estimation Under Corruption
Lemma C.1.
Let $P_{t}(a)$ represent the likelihood of action $a_{t}$ being preferred over $a$, where $P_{t}(a):=\sigma(\mu(a_{t})-\mu(a))$. Likewise, the corrupted version is defined as $\hat{P}_{t}(a)=\sigma(\mu(a_{t})-\mu(a)+c_{t}(a_{t},a))$. The smoothed version of $P_{t}$ over $\mathcal{A}$ is $\bar{P}_{t}(a):=\mathbb{E}_{x\in\mathbb{B}}\left[P_{t}(\mathcal{P}_{\mathcal{A}}(a+\delta x))\right]$. Let $a^{\prime}_{t}=\mathcal{P}_{\mathcal{A}}(a_{t}+\delta u_{t})$, where $u_{t}$ is uniformly sampled from $\mathbb{S}$ and $\hat{g}_{t}=-\frac{d}{\delta}\hat{\mathcal{F}}\left(a^{\prime}_{t},a_{t}\right)u_{t}$. Let $b_{t}:=\frac{d}{\delta}\left(\hat{P}_{t}(a^{\prime}_{t})-P_{t}(a^{\prime}_{t})\right)u_{t}$. We have $\mathbb{E}(\hat{g}_{t}|a_{t})=\mathbb{E}(\nabla\bar{P}_{t}(a_{t})|a_{t})-\mathbb{E}(b_{t}|a_{t}).$
Proof  (Proof for Lemma C.1). Since $a^{\prime}_{t}:=\mathcal{P}_{\mathcal{A}}(a_{t}+\delta u_{t})$, we have
$\mathbb{E}(\hat{g}_{t}|a_{t})$ $=\mathbb{E}_{u_{t}}(\mathbb{E}(\hat{g}_{t}|a_{t},u_{t}))$
$=-\frac{d}{\delta}\mathbb{E}_{u_{t}}\left(\mathbb{E}(\hat{\mathcal{F}}(a^{\prime}_{t},a_{t})u_{t}|a_{t},u_{t})\right)$
$=-\frac{d}{\delta}\mathbb{E}\left(\hat{P}_{t}(a^{\prime}_{t})u_{t}|a_{t}\right)$
$=-\frac{d}{\delta}\mathbb{E}\left(\left[P_{t}(a^{\prime}_{t})+\hat{P}_{t}(a^{\prime}_{t})-P_{t}(a^{\prime}_{t})\right]u_{t}|a_{t}\right)$
$=\nabla\mathbb{E}_{x\in\mathbb{B}}\left(P_{t}(\mathcal{P}_{\mathcal{A}}(a_{t}+\delta x))|a_{t}\right)-\frac{d}{\delta}\mathbb{E}\left[\left(\hat{P}_{t}(a^{\prime}_{t})-P_{t}(a^{\prime}_{t})\right)u_{t}|a_{t}\right]$ (14)
$=\nabla\bar{P}_{t}(a_{t})-\frac{d}{\delta}\mathbb{E}\left[\left(\hat{P}_{t}(a^{\prime}_{t})-P_{t}(a^{\prime}_{t})\right)u_{t}|a_{t}\right]$
$=\mathbb{E}(g_{t}|a_{t})-\frac{d}{\delta}\mathbb{E}\left[\left(\hat{P}_{t}(a^{\prime}_{t})-P_{t}(a^{\prime}_{t})\right)u_{t}|a_{t}\right].$ (15)
We get the equation (14) by using Lemma C.2. We get equation (15) by using Lemma C.3.
Lemma C.2 (Lemma 2 in Yue and Joachims (2009)).
Fix $\delta>0$, over random unit vector $u$, we have
$\mathbb{E}[P_{t}(\mathcal{P}_{\mathcal{A}}(a+\delta u))u]=\frac{\delta}{d}\nabla\bar{P}_{t}(a).$
Lemma C.3 (Lemma 1 in Yue and Joachims (2009)).
$\mathbb{E}_{\mathcal{F},u}[\mathcal{F}(a^{\prime}_{t},a_{t})u]=-\mathbb{E}_{u}[P_{t}(a^{\prime}_{t})u]$.
C.1.2 Proof for Lemma 2: Regret Decomposition
Lemma (Lemma 2 restated).
Choose $\lambda=\frac{L_{\sigma}}{L_{\sigma}-\delta L_{\mu}L_{\sigma}},\gamma=R/\sqrt{T}$, for $b_{t}$ defined in Lemma C.1, we have
[…]
Proof  (Proof for Lemma 2). Because of Lemma C.4, we have
$\mathbb{E}\left[\sum^{T}_{t=1}\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*})\right]\leq\sum^{T}_{t=1}\mathbb{E}\left(\lambda\nabla\bar{P}_{t}(a_{t})(a_{t}-a^{*})+(3+\lambda)\delta L\right)$
$=\lambda\sum^{T}_{t=1}\mathbb{E}(\mathbb{E}(g_{t}|a_{t})^{\top}(a_{t}-a^{*}))+(3+\lambda)\delta LT$
$=\lambda\sum^{T}_{t=1}\mathbb{E}(\mathbb{E}(\hat{g}_{t}+b_{t}|a_{t})^{\top}(a_{t}-a^{*}))+(3+\lambda)\delta LT$
$=\lambda\sum^{T}_{t=1}\mathbb{E}(\hat{g}_{t}^{\top}(a_{t}-a^{*}))+\lambda\sum^{T}_{t=1}\mathbb{E}(b_{t}^{\top}(a_{t}-a^{*}))+(3+\lambda)\delta LT$
Lemma C.4 (Lemma 4 in Yue and Joachims (2009)).
Fix $\delta\in(0,\frac{L_{\sigma}}{L_{\nu}L_{\sigma}})$ and define $\lambda=\frac{L_{\sigma}}{L_{\sigma}-\delta L_{\nu}L_{2}}$, we have
$\mathbb{E}\left[\sum^{T}_{t=1}\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*})\right]\leq\sum^{T}_{t=1}\mathbb{E}\left(\lambda\nabla\bar{P}_{t}(a_{t})^{\top}(a_{t}-a^{*})+(3+\lambda)\delta L\right).$
Since $a_{t+1}=\mathcal{P}_{\mathcal{A}}(a_{t}-\eta\hat{g}_{t})$, and $\eta=\frac{\gamma\delta}{d}$, by applying the telescoping sum and because $a_{1}=0$, $\|g_{t}\|_{2}\leq G,\forall t$, we have
$\lambda\sum^{T}_{t=1}\mathbb{E}\left(\hat{g}_{t}^{\top}(a_{t}-a^{*})\right)\leq\lambda\left(\frac{R^{2}}{2\eta}+\frac{T\eta G^{2}}{2}\right).$
Noticing that $\|b_{t}\|_{2}\leq\frac{d}{\delta}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)$. Moreover, $\|a_{t}-a^{*}\|_{2}\leq 2R$. Therefore, we can control the observation error by the following.
C.1.3 Proof for Lemma C.5: Observation Error
Lemma C.5.
$\sum^{T}_{t=1}\mathbb{E}\left(b_{t}^{\top}(a_{t}-a^{*})\right)\leq 2RdL_{\sigma}C/\delta$.
Proof  (Proof for Lemma C.5).
$\sum^{T}_{t=1}\mathbb{E}(b_{t}(a_{t}-a^{*}))$ $\leq 2R\frac{d}{\delta}\sum^{T}_{t=1}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)$
$\leq 2R\frac{dL_{\sigma}}{\delta}\sum^{T}_{t=1}|c_{t}(a_{t},a^{\prime}_{t})|$
$=2R\frac{dL_{\sigma}C}{\delta}.$
Lemma C.6 (Lemma 3 in Yue and Joachims (2009)).
$\text{Reg}_{T}\leq 2\mathbb{E}\left[\sum^{T}_{t=1}\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*})\right]+5\delta LT$
Set $G=\frac{d}{\delta}$, and by Lemma C.6, we have
$\text{Reg}_{T}\leq\lambda\left(\frac{2Rd\sqrt{T}}{\delta}+13\delta LT\right)+4R\frac{\lambda dL_{\sigma}C}{\delta}.$
Therefore, by setting $\delta=\frac{\sqrt{2Rd}}{\sqrt{13L}T^{\alpha}},\gamma=\frac{R}{\sqrt{T}}$, and $T>\left(\frac{\sqrt{2Rd}L_{v}L_{2}}{\sqrt{13L}L_{\sigma}}\right)^{4}$, we have
$\text{Reg}_{T}\leq 2\lambda_{T}\sqrt{26RdL}T^{1-\alpha}+2L_{\sigma}\sqrt{26RdL}T^{\alpha}C.$
$\lambda_{T}=\frac{L_{\sigma}\sqrt{13L}T^{\alpha}}{L_{\sigma}\sqrt{13L}T^{\alpha}-L_{v}L_{2}\sqrt{2Rd}}$.
C.2 Proof for Corollary 1: Minimax Lower Bound for DBGD
Theorem 2 implies that DBGD could afford agnostic corruption level $O(T^{3/4})$. This implies that there exists a hard instance which makes DBGD attain regret in order of $\Omega(T^{3/4})$, formally described as follows.
Corollary (Corollary 1 restated).
There exists a linear utility function $\mu$ and a link function $\sigma$ such that choosing $\alpha=1/4$, $\text{Reg}_{T}$ suffered by Algorithm 1 is at least $\Omega\left(T^{3/4}\right)$.
Proof  (Proof for Corollary 1). To start a proof by contradiction, we assume that Theorem 2 is false. Specifically, for all problem instance $\mu,\sigma$, there exist a $\epsilon>0$ such that $\text{Reg}_{T}\leq O(T^{3/4-\epsilon})$. We assume $\epsilon$ to be the least possible. In particular, there exists a pair of instance $\mu,\sigma$ such that $\text{Reg}_{T}=c_{0}T^{3/4-\epsilon}$, $c_{0}$ is a positive constant. In other words, it says that there exists a pair of instance $\mu,\sigma$ such that DBGD suffers regret in $\Theta(T^{3/4-\epsilon})$.
Consider the following problem instance. $\mu(a)=\theta^{\top}a$, specifically $\mu$ is linear function. Let $d=2$ with action set $\mathcal{A}_{1}:=\{(a_{1},a_{2}):a_{1}\geq 0,a_{2}\geq 0,\frac{1}{2}a_{1}+a_{2}-\frac{1}{4}\leq 0\}$ with $\theta=[\frac{1}{2},\frac{1}{2}]$. The optimal arm is $a_{1}=[\frac{1}{2},0]$, which is unique. Let the link function $\sigma(x)=\frac{1}{2}+\frac{1}{2}x$. It is easy to verify that it is rotational symmetric and Lipschitz. Denote $R_{T}$ as the regret occurred by DBGD on this problem instance. We know $R_{T}\leq O(T^{\frac{3}{4}-\epsilon})$. Consequently, it implies that $a_{2}$ is at most proposed $8c_{0}T^{\frac{3}{4}-\epsilon}$ times. This is because $R_{T}\leq c_{0}T^{\frac{3}{4}-\epsilon}$ and if the proposed action pair $(a_{t},a^{\prime}_{t})$ including $a_{2}$ at round $t$, it occurs regret at least $\frac{1}{8}$.
Now consider a different problem instance with the same $\mu,\sigma$ but different action set, which is $\mathcal{A}_{2}:=\{(a_{1},a_{2}):a_{1}\geq 0,a_{2}\geq 0,\frac{3}{2}a_{1}+a_{2}-\frac{3}{4}\leq 0\}$. The optimal arm in this action set $a_{2}=[0,\frac{3}{4}]$, which is also unique. Proposing arm $a_{1}=[\frac{1}{2},0]$ occurs at least $\frac{1}{8}$ regret. Consider an adversary that pulls the utility of $a_{2}$ from $\frac{3}{8}$ down to $\frac{1}{8}$ whenever it is pulled at a cost of $c_{t}(A_{t})=\frac{1}{4}$. It is equivalent to say that every time when the proposed arm $w$ falls in the region $\mathcal{W}=\{a_{1}\geq 0,a_{2}\geq 0,\frac{1}{2}a_{1}+a_{2}-\frac{1}{4}\geq 0,\frac{3}{2}a_{1}+a_{2}-\frac{3}{4}\leq 0\}$, we assume that the utility is sampled from $\theta^{\top}\mathcal{P}_{\mathcal{A}_{1}}(w)$ and $|c_{t}(w)|=|\theta^{\top}\mathcal{P}_{\mathcal{A}_{1}}(w)-\theta^{\top}w|\leq\frac{1}{4}$. Choosing $C=2c_{0}T^{\frac{3}{4}-\epsilon}$, the adversary can afford to corrupt $8c_{0}T^{\frac{3}{4}-\epsilon}$ times.
This means that as long as the adversary is corrupting, the utility observed in the first problem instance is exactly the same as the utility observed in the second problem instance, in which we pull $a_{2}$ as most $8c_{0}T^{\frac{3}{4}-\epsilon}$. However, $a_{2}$ is the optimal arm in the second problem instance, which implies that the regret occurred on the second problem instance is at least $T-8c_{0}T^{\frac{3}{4}-\epsilon}$, which is linear in $T$. This contradicts Theorem 2, which says that when agnostic corruption $C=2c_{0}T^{\frac{3}{4}-\epsilon}$, the regret upper bound is $O(T^{1-\epsilon})$, which is sublinear. To reconcile the conflicts, we must have Proposition 1 true to make the agnostic corruption $C\geq\Omega(T^{\frac{3}{4}})$ to make the parallel world argument valid.
C.3 Proof for Matching Regret Upper Bound in Theorem 1
In this section, we consider the setting when human has strongly concave utility function. When $\rho$ is known, we optimally adjusts the learning rate of the algorithm, Noisy Comparision-based Stochastic Mirror Descent (NC-SMD), proposed in Kumagai (2017), to achieve a balance between optimizing the learning objective (1) and controlling the gradient bias arising from the corrupted dueling feedback. Then we show that it attains the matching regret upper bound of $\tilde{O}(d\max\{\sqrt{T},T^{\rho}\})$ under corruption induced by $\rho$-imperfect user.
For the ease of delivering proof, in this section, we work with the cost function $f$, opposite of the utility function $\mu$, We restate the modelling of inaccuracies from human feedback and the learning objective using $f$ as follows.
Assumption C.1.
(Corrupted Utility-Based Dueling Bandit). For a pair of actions $(a,a^{\prime})\in\mathcal{A}\times\mathcal{A}$ and the corruption level $c(a,a^{\prime})$, there exists function $f:\mathcal{A}\rightarrow\mathbb{R}$ and $\sigma:\mathbb{R}\rightarrow[0,1]$ such that the probability that $a\succ a^{\prime}$ happens of receiving corrupted dueling feedback can be represented follows
^$\hat{\mathbb{P}}(a\succ a^{\prime})=\sigma(f(a^{\prime})-f(a)+c(a,a)).$
Additionally, we assume that the cost function $f$, the link function $\sigma$, and the action space $\mathcal{A}$ exhibit the following properties, which align precisely with those assumed in Kumagai (2017).
Assumption C.2.
The cost function $f:\mathcal{A}\rightarrow\mathbb{R}$ is twice continuously differentiable, $L$-Lipschitz, $\alpha$-strongly convex and $\beta$-smooth with respect to the Euclidean norm.
Definition C.1.
(Strong Convexity). A function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $\alpha$-strongly convex over the set $\mathcal{A}\subset\mathbb{R}^{d}$ if for all $x,y\in\mathcal{A}$ it holds that
$f(y)\geq f(x)+\nabla f(x)^{\top}(y-x)+\frac{\alpha}{2}\|y-x\|^{2}_{2}.$
Definition C.2.
(Smoothness). A function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $\beta$-smooth over the set $\mathcal{A}\subset\mathbb{R}^{d}$ if for all $x,y\in\mathcal{A}$ it holds that
$f(y)\leq f(x)+\nabla f(x)^{\top}(y-x)+\frac{\beta}{2}\|y-x\|^{2}_{2}.$
Assumption C.3.
Let $B:=\sup_{a,a^{\prime}\in\mathcal{A}}f(a^{\prime})-f(a)$, $f$ is defined in Assumption C.2. The link function $\sigma:\mathbb{R}\rightarrow[0,1]$ is three times differentiable and rotational-symmetric. Its first derivative is positive and monotonically non-increasing on $[0,B]$.
Assumption C.4.
We assume the action space $\mathcal{A}\in\mathbb{R}^{d}$ is convex, compact and has non-empty interior.
Definition C.3.
(Self-concordance.) A function $\mathcal{R}:\text{int}(\mathcal{A})\rightarrow\mathbb{R}$ is self-concordant if it satisfies
  1. 1.
    $\mathcal{R}$ is three times continuously differentiable, convex, and approaches infinity along any sequence of points approaching the boundary of $\text{int}(\mathcal{A})$.
  2. 2.
    For every $h\in\mathbb{R}^{d}$ and $x\in\text{int}(\mathcal{A})$, $|\nabla^{3}\mathcal{R}(x)[h,h,h]|\leq 2\left(h^{\top}\nabla^{2}\mathcal{R}(x)h\right)^{\frac{3}{2}}$ holds, where $|\nabla^{3}\mathcal{R}(x)[h,h,h]:=\frac{\partial^{3}\mathcal{R}}{\partial t_{1}\partial t_{2}\partial t_{3}}(x+t_{1}h+t_{2}h+t_{3}h)|_{t_{1}=t_{2}=t_{3}=0}.$
In addition to these two conditions, if for every $h\in\mathbb{R}^{d}$ and $x\in\text{int}(\mathcal{A})$, $|\nabla\mathcal{R}(x)^{\top}h\leq\nu^{\frac{1}{2}}(h^{\top}\nabla^{2}\mathcal{R}(x)h)^{\frac{1}{2}}|$ for a positive real number $\nu$, $\mathcal{R}$ is $\nu$-self-concordant.
Through interactions with the user, the learner $\mathcal{L}$ endeavors to minimize the dueling regret, as defined below.
$\text{Reg}^{\text{DB}}_{T}=\mathbb{E}\left(\sum^{T}_{t=1}\sigma(f(a_{t})-f(a^{*}))+\sigma(f(a^{\prime}_{t})-f(a^{*}))-1\right).$ (16)
Our minimization objective in (16) aligns with the regret definition in Yue and Joachims (2009). Under Assumptions C.1, C.2, and C.3, the regret defined in (16) is equivalent to the regret defined in Kumagai (2017). Assumption C.2 suggests the existence of a unique minimizer $a^{*}$ for the cost function $f$ due to its strict convexity. Assumptions C.2 and C.4 imply that the diameter $R:=\sup_{a,a^{\prime}\in\mathcal{A}}|a-a^{\prime}|$ and $B$ are finite. Assumption C.3 implies the existence of positive constants $l_{0}$, $L_{0}$, $B_{2}$, and $L_{2}$ such that the first derivative $\sigma^{\prime}$ of the link function is bounded, satisfying $l_{0}\leq\sigma^{\prime}\leq L_{0}$ on $[-B,B]$. Furthermore, the second derivative $\sigma^{{}^{\prime\prime}}$ is bounded above by $B_{2}$ and is $L_{2}$-Lipschitz on $[-B,B]$. Lemma C.7 establishes the relationship between $\text{Reg}^{\text{FO}}_{T}$ and $\text{Reg}^{\text{DB}}_{T}$.
Lemma C.7 (Lemma 12 in Kumagai (2017)).
$\frac{\text{Reg}^{\text{DB}}_{T}}{L_{0}}\leq\text{Reg}^{\text{FO}}_{T}\leq\frac{\text{Reg}^{\text{DB}}_{T}}{l_{0}}.$ (17)
C.3.1 Noisy Comparison-based Stochastic Mirror Descent
Input: Learning rate $\eta_{\rho}$, $\nu$-self-concordant function $\mathcal{R}$, time horizon $T$, tuning parameters $\lambda$, $\mu$.
1 Initialize $a_{1}=\arg\min_{a\in A}\mathcal{R}(a)$
2 for $t\in[T]$ do
3       Update $\mathcal{R}_{t}(a)=\mathcal{R}(a)+\frac{\lambda\eta_{\rho}}{2}\sum^{t}_{i=1}\|a-a_{i}\|^{2}+\mu\|a\|^{2}$
4       Pick a unit vector $u_{t}$ uniformly at random
5       Receive the corrupted feedback ^$\hat{\mathcal{F}}(a^{\prime}_{t},a_{t})$, for $a_{t}$ and $a^{\prime}_{t}=a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-1/2}u_{t}$
6       Compute the corrupted gradient: $\hat{g}_{t}=\hat{\mathcal{F}}(a^{\prime}_{t},a_{t})d\nabla^{2}\mathcal{R}_{t}(a_{t})^{1/2}u_{t}$
7       Set $a_{t+1}=\nabla\mathcal{R}_{t}^{-1}(\nabla\mathcal{R}_{t}(a_{t})-\eta_{\rho}\hat{g}_{t})$
Output : $a_{T+1}$
Algorithm 3 Noisy Comparison-based Stochastic Mirror Descent (NC-SMD)
Remark C.1.
Note that $a^{\prime}_{t}=a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}x$ for $x\in\mathbb{B}$ is included in $\mathcal{A}$ due to the properties of the Dikin ellipsoid.
Now we begin to prove the matching regret upper bound in Theorem 1.
Proposition C.1 (Full Statement of Matching Regret Upper Bound in Theorem 1).
Let the number of rounds $T$ be fixed and $T>\sqrt{2LL_{0}R}$. Consider a corruption induced by $\rho$-imperfect user (Def. 1) with known $\rho$. Under Assumptions C.2, C.3, and C.4, choose $\lambda\leq l_{0}\alpha/2$, $\mu\geq(L^{3}_{0}L_{2}/\lambda)^{2}$, and set the learning rate $\eta_{\rho}:=\frac{\sqrt{\log T}}{dT^{\max\{1/2,\rho\}}}$ for Algorithm 3. We have we have:
$\text{Reg}^{\text{DB}}_{T}\leq O\left(\max\left\{d\sqrt{T\log T},dT^{\rho}\log T\right\}\right).$
Proof  (Proof for Proposition C.1). The formal proof of Proposition C.1 relies on Lemmas C.8, C.9, and C.10. Lemma C.8 establishes a connection between the gradient estimation corrupted by the bias term $b_{t}$ and the quantity $\bar{P}_{t}(a)$ (defined in Lemma C.8), which is the main component we aim to analyze in the regret definition (16). Lemma C.9 decomposes the regret into two parts: a regret of decision independent of $b_{t}$ and an error in observation summarizing the cumulative impact of $b_{t}$.
The optimization error can be explicitly upper bounded by $T$ and $\eta_{\rho}$ using techniques similar to those employed in Kumagai (2017). The upper bound of the estimation error is further established in Lemma C.10, wherein we utilize the $\alpha$-strong convexity of $f$ to establish a connection between the convergence of the $L_{2}$-norm of the action $\|a_{t}-a^{*}\|_{2}$ to itself.
Lemma C.8 (Gradient Estimation under Corruption).
Let $P_{t}(a)$ represent the likelihood of action $a_{t}$ being preferred over $a$, where $P_{t}(a):=\sigma(f(a)-f(a_{t}))$. Likewise, the corrupted version is defined as $\hat{P}_{t}(a)=\sigma(f(a)-f(a_{t})+c_{t}(a_{t},a^{\prime}_{t}))$. The smoothed version of $P_{t}$ over $\text{int}(\mathcal{A})$ is $\bar{P}_{t}(a):=\mathbb{E}_{x\in\mathbb{B}}\left[P_{t}(a+\nabla^{2}R_{t}(a_{t})^{-\frac{1}{2}}x)\right]$. $\mathcal{R}_{t}(a)$ is defined in Algorithm 3. Let $a^{\prime}_{t}=a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t}$, where $u_{t}$ is uniformly sampled from $\mathbb{S}$. Let $b_{t}:=d\left(P_{t}(a^{\prime}_{t})-\hat{P}_{t}(a^{\prime}_{t})\right)\nabla^{2}\mathcal{R}_{t}(a_{t})^{\frac{1}{2}}u_{t}$, we have
$\mathbb{E}\left(\hat{g}_{t}|a_{t}\right)=\nabla\bar{P}_{t}(a_{t})-\mathbb{E}\left(b_{t}|a_{t}\right).$
Lemma C.9 (Regret Decomposition).
Define $C_{0}:=(2\nu+4L_{0}\beta+4B_{2}L^{2}+LL_{0}\beta)/\lambda$, and let $a^{*}_{T}=\frac{1}{T}a_{1}+(1-\frac{1}{T})a^{*}$. By Assumption C.3, executing Algorithm 3, we obtain:
[…]
Lemma C.10 (Estimation Error Upper Bound).
Under Assumption C.2, and C.4, we have
$\mathbb{E}\left\{\sum^{T}_{t=1}b^{\top}_{t}(a_{t}-a^{*}_{T})\right\}$ $\leq d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}\right\}+C_{2}dC+2d\sqrt{\lambda\eta_{\rho}T},$
where we have $\lambda^{*}_{\mathcal{R}}:=\sup_{a\in\mathcal{A}}\lambda_{\max}\left(\nabla^{2}\mathcal{R}(a)\right)$, and $m_{t},n_{t},C_{2}$ defined as
$m_{t}:=\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right),n_{t}:=\min\left(2R,\sqrt{\frac{2}{\alpha}\left(f(a_{t})-f(a^{*})\right)}\right),C_{2}:=2L_{\sigma}R\sqrt{\lambda^{*}_{R}+2\mu}.$
Now, we present the proof of Proposition C.1. The key technical innovation lies in realizing that the upper bound of $\text{Reg}^{\text{DB}}_{T}$ contains itself. This implies that we can initially obtain a rough upper bound and then refine it until it becomes tight. The proof follows a mathematical induction argument. For the ease of notation, we reparameterize $C=C_{\kappa}T^{\rho}$. In the following proof, we consider $\rho>0.5$ to make $T^{\rho}=\max\{\sqrt{T},T^{\rho}\}$. If $\rho\leq 0.5$, the following proof can be easily modified to make upper bounds involving $T^{\rho}$ as $\sqrt{T}$.
In the first step, we employ a rough upper bound, specifically $\mathbb{E}\left(n_{t}\right)\leq 2R$. This yields $\text{Reg}^{DB}_{T}\leq O(dT^{\frac{\rho+1}{2}})$. Since $\text{Reg}^{DB}_{T}$ and $\text{Reg}^{\text{FO}}_{T}$ share the same order (see Equation (17)), it implies that $\sum^{T}_{t=1}\mathbb{E}\left(n_{t}\right)$ must be sublinear as well. Therefore, in the second step, we can use Abel’s Summation Equation (Lemma C.16) to refine the upper bound of $\mathbb{E}\left(\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}\right)$.
In particular, we rigorously prove the following claim:
The Induction Claim: Given the number of iteration $T$. $T>\sqrt{2LL_{0}R}$. At k-th step, we have $\text{Reg}^{\text{DB}}_{T}\leq 144dK\sqrt{\log T}T^{\rho+\frac{\frac{3}{2}-\rho}{2^{k+1}}}$, where $K:=\max\left\{C_{0},\frac{C_{2}C_{\kappa}}{\sqrt{\log T}},L^{2}_{\sigma}C^{2}_{\kappa},2\sqrt{\lambda}RC_{\kappa}L_{\sigma}\right\}$.
The Base Case: When $k=1$, let $C_{3}=C_{2}dC+2d\sqrt{\lambda\eta_{\rho}T}$, we have
$\text{Reg}^{\text{DB}}_{T}$ $\leq\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2LL_{0}R+2C_{3}+2d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left(\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}\right).$ (18)
$\leq\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2LL_{0}R+2C_{3}+4dR\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}m_{t}\right\}$ (19)
$\leq\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2LL_{0}R+2C_{3}+4dR\sqrt{\lambda\eta_{\rho}T}\left(\sum^{T}_{t=1}m_{t}\right)$ (20)
$\leq\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2LL_{0}R+2C_{3}+4dL_{\sigma}R\sqrt{\lambda\eta_{\rho}}C_{\kappa}T^{\rho}.$ (21)
We get Equation (18) according to Lemma C.9 and C.10. We get Equation (19) because of $\mathbb{E}\left(n_{t}\right)\leq 2R$. We get Equation (20) by Abel’s Summation Equation (see Lemma C.16). We get get Equation (21) because of corruption budget. Choosing the learning rate $\eta_{\rho}$ according to Proposition C.1, we obtain $\text{Reg}^{\text{DB}}_{T}\leq 144\sqrt{\log T}KdT^{\frac{\rho+1}{2}}$. This confirms the validity of the claim when $k=1$.
The Induction Argument: Let’s assume that the claim holds true for a general step $k$. In the following, we will demonstrate that the claim also holds true for step $k+1$. Because of Equation (17) and the induction claim, we obtain $\text{Reg}^{\text{FO}}_{T}\leq\frac{144}{l_{0}}dK\sqrt{\log T}T^{\rho+\frac{3/2-\rho}{2^{k+1}}}.$ This suggests that
$\sum^{T}_{t=1}\mathbb{E}\left(f(a_{t})-f(a^{*})\right)\leq\frac{144}{l_{0}}dK\sqrt{\log T}T^{\rho+\frac{3/2-\rho}{2^{k+1}}}.$
By Jensen’s inequality, we have
$\mathbb{E}(n_{t})=\min\left\{2,\sqrt{\frac{2}{\alpha}}\mathbb{E}\left(\sqrt{f(a_{t})-f(a^{*})}\right)\right\}\leq\sqrt{\frac{2}{\alpha}}\sqrt{\mathbb{E}\left(f(a_{t})-f(a^{*})\right)}.$
Therefore we have
$\mathbb{E}\left(\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}\right)\leq\sqrt{\frac{2}{\alpha}}\sum^{T}_{t=1}\sqrt{t}m_{t}\sqrt{\mathbb{E}\left(f(a_{t})-f(a^{*})\right)}.$
By the definition of monotone adversary, we have $|c_{t}(a_{t},a^{\prime}_{t})|$ is upper bounded by $c_{k}t^{-1+\rho}$.
Lemma C.11.
Consider $\sum^{T}_{t=1}n^{2}_{t}\leq C^{\prime}T^{\frac{1}{2}+\alpha}$, with constants $\alpha\geq 0$, $C^{\prime}>0$. In addition, $n_{t}$ is uniformly upper bounded by a constant $K$. Specifically, we have $0\leq n_{t}\leq K,\forall t$. $m_{t}\leq c_{k}t^{\rho-1}$ with constant $c_{k}\geq 0$. Then we have $\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}\leq 5\sqrt{C^{\prime}}c_{k}T^{-\frac{1}{4}+\frac{\alpha}{2}+\rho}$.
Therefore, using Lemma C.11 (details of proof is listed at Appendix C.3.5), we can upper bound $\sqrt{\lambda}\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}$ by
$\sqrt{\lambda}\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}\leq 60L_{\sigma}\sqrt{\frac{2\lambda}{\alpha l_{0}}}(\log T)^{0.25}C_{k}\sqrt{dK}T^{\frac{3}{2}\rho+\frac{3/2-\rho}{2^{k+2}}}$
Since $\lambda\leq\alpha l_{0}/2$, choosing $\eta_{\rho}=\frac{1}{2d}\frac{\sqrt{\log T}}{T^{\rho}}$, we have
$\text{Reg}^{\text{DB}}_{T}$ $\leq 2d\left(C_{0}+\frac{C_{2}}{\sqrt{\log T}}+60C_{\kappa}\sqrt{K}T^{\frac{3/2-\rho}{2^{k+2}}}\right)\sqrt{\log T}T^{\rho}+12Kd\sqrt{\log T}T^{\rho}$
$\leq 144dK\sqrt{\log T}T^{\rho+\frac{3/2-\rho}{2^{k+2}}}.$
The claim holds true for step $k+1$. Therefore, it implies that the claim holds true for all k. Repeating this process infinitely many times, we obtain
$\text{Reg}^{\text{DB}}_{T}\leq O\left(d\sqrt{\log T}T^{\rho}\right),$
which completes the proof when $\rho\in[0.5,1]$.
C.3.2 Proof for Lemma C.8: Gradient Estimation Under Corruption
Proof  (Proof for Lemma C.8). If we do not have corruption (i.e. the probability for observing noisy comparative feedback $\mathcal{F}(a,a^{\prime})=1$ is based on true cost difference $f(a^{\prime})-f(a)$), then the uncorrupted gradient $g_{t}:=\mathcal{F}(a^{\prime}_{t},a_{t})d\nabla^{2}\mathcal{R}_{t}(a_{t})^{1/2}u_{t}$ should be an unbiased estimate of $\nabla\bar{P}_{t}(a_{t})$, i.e.
$\mathbb{E}(g_{t}|a_{t})=\nabla\bar{P}_{t}(a_{t}).$
The proof is similar to Lemma 2.1 in Flaxman et al. (2005). Using the Law of total expectation, we have
$\mathbb{E}(g_{t}|a_{t})$ $=\mathbb{E}_{u_{t}}(\mathbb{E}(g_{t}|a_{t},u_{t}))$
$=\mathbb{E}_{u_{t}}\left(d\mathbb{E}(P_{t}(a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t})\nabla^{2}\mathcal{R}_{t}(a_{t})^{\frac{1}{2}}u_{t}|a_{t},u_{t})\right)$
$=d\mathbb{E}(P_{t}(a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t})\nabla^{2}\mathcal{R}_{t}(a_{t})^{\frac{1}{2}}u_{t}|a_{t})$
$=\nabla\mathbb{E}_{x\in\mathbb{B}}\left(P_{t}(a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}x|a_{t})\right)$
$=\nabla\bar{P}_{t}(a_{t}).$
We get the second inequality by using the definition of $\mathcal{F}(a^{\prime}_{t},a_{t})$. We use the Stroke’s Theorem to get the second last equality. The gradient which we get to perform gradient descent $\hat{g}_{t}$ is corrupted. If we let $a^{\prime}_{t}=a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t}$, we have
$\mathbb{E}(\hat{g}_{t}|a_{t})$ $=\mathbb{E}_{u_{t}}(\mathbb{E}(\hat{g}_{t}|a_{t},u_{t}))$
$=\mathbb{E}_{u_{t}}\left(d\mathbb{E}(\hat{P}_{t}(a^{\prime}_{t})\nabla^{2}\mathcal{R}_{t}(a_{t})^{\frac{1}{2}}u_{t}|a_{t},u_{t})\right)$
$=d\mathbb{E}\left(\hat{P}_{t}(a^{\prime}_{t})\nabla^{2}\mathcal{R}_{t}(a_{t})^{\frac{1}{2}}u_{t}|a_{t}\right)$
$=d\mathbb{E}\left(\left[P_{t}(a^{\prime}_{t})+\hat{P}_{t}(a^{\prime}_{t})-P_{t}(a^{\prime}_{t})\right]\nabla^{2}R_{t}(a_{t})^{\frac{1}{2}}u_{t}|a_{t}\right)$
$=\nabla\mathbb{E}_{x\in\mathbb{B}}\left(P_{t}(a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}x|a_{t})\right)+d\mathbb{E}\left[\left(\hat{P}_{t}(a^{\prime}_{t})-P_{t}(a^{\prime}_{t})\right)\nabla^{2}\mathcal{R}_{t}(a_{t})^{\frac{1}{2}}u_{t}|a_{t}\right]$
$=\nabla\bar{P}_{t}(a_{t})+d\mathbb{E}\left[\left(\hat{P}_{t}(a^{\prime}_{t})-P_{t}(a^{\prime}_{t})\right)\nabla^{2}\mathcal{R}_{t}(a_{t})^{\frac{1}{2}}u_{t}|a_{t}\right]$
$=\mathbb{E}(g_{t}|a_{t})+d\mathbb{E}\left[\left(\hat{P}_{t}(a^{\prime}_{t})-P_{t}(a^{\prime}_{t})\right)\nabla^{2}\mathcal{R}_{t}(a_{t})^{\frac{1}{2}}u_{t}|a_{t}\right].$
We get the second last equality by using the fact that $g_{t}$ is unbiased, which we proved earlier. If we defined $b_{t}$ as
$b_{t}:=d\left(P_{t}(a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t})-\hat{P}_{t}(a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t})\right)\nabla^{2}\mathcal{R}_{t}(a_{t})^{\frac{1}{2}}u_{t},$
we get
$\mathbb{E}(g_{t}|a_{t})=\mathbb{E}(\hat{g}_{t}|a_{t})+\mathbb{E}\left[b_{t}|a_{t}\right].$
C.3.3 Proof for Lemma C.9: Regret Decomposition
Proof  (Proof for Lemma C.9). The cornerstone of the analysis below is to separate the impact of $b_{t}$ from the total regret.
$\text{Reg}^{\text{DB}}_{T}\leq 2\mathbb{E}\left[\sum^{T}_{t=1}(P_{t}(a_{t})-P_{t}(a^{*}_{T})\right]+\frac{LL_{0}\beta}{\lambda\eta}+2LL_{0}R$
$\leq 2\left(\mathbb{E}\left\{\sum^{T}_{t=1}(\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*}_{T}))\right\}+\mathbb{E}\left\{\sum^{T}_{t=1}(P_{t}(a_{t})-\bar{P}_{t}(a_{t}))\right\}+\mathbb{E}\left\{\sum^{T}_{t=1}(\bar{P}_{t}(a^{*}_{T})-P_{t}(a^{*}_{T}))\right\}\right)$
$+\frac{LL_{0}\beta}{\lambda\eta}+2LL_{0}R$
$\leq 2\mathbb{E}\left\{\sum^{T}_{t=1}(\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*}_{T}))\right\}+\frac{4L_{0}\beta+4B_{2}L^{2}+LL_{0}\beta}{\lambda\eta_{\rho}}\log T+2LL_{0}R.$
We get the first inequality by using Lemma C.12. We get the last inequality by because
$\bar{P}_{t}(a)-P_{t}(a)\leq\frac{L_{0}\beta+B_{2}L^{2}}{2}\|\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t}\|^{2}\leq\frac{L_{0}\beta+B_{2}L^{2}}{\lambda\eta_{\rho}t},$
and
$\mathbb{E}\left\{\sum^{T}_{t=1}(P_{t}(a_{t})-P_{t}(a^{*}_{T}))\right\}$ $\leq\mathbb{E}\left\{\sum^{T}_{t=1}(\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*}_{T}))\right\}+2\frac{L_{0}\beta+B_{2}L^{2}}{\lambda\eta_{\rho}}\log T.$
Lemma C.12 (Lemma 5 in Kumagai (2017)).
$\text{Reg}^{\text{DB}}_{T}\leq 2\mathbb{E}\left[\sum^{T}_{t=1}(P_{t}(a_{t})-P_{t}(a^{*}_{T})\right]+\frac{LL_{0}\beta}{\lambda\eta}+2LL_{0}R.$
Then it remains to bound $\mathbb{E}\{\sum^{T}_{t=1}\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*}_{T})\}$. And we have the following.
$\mathbb{E}\{\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*}_{T})\}$ $\leq\mathbb{E}\left[\nabla\hat{P}^{\top}_{t}(a_{t})(a_{t}-a^{*}_{T})-\frac{l_{0}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]$
$=\mathbb{E}\left[g_{t}^{\top}(a_{t}-a^{*}_{T})-\frac{l_{0}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]$
$=\mathbb{E}\left[(\hat{g}_{t}+b_{t})^{\top}(a_{t}-a^{*}_{T})-\frac{l_{0}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]$
$=\mathbb{E}\left[\hat{g}_{t}^{\top}(a_{t}-a^{*}_{T})-\frac{l_{0}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]+\mathbb{E}\{b^{\top}_{t}(a_{t}-a^{*}_{T})\},$
where the first and second equality is resulted from Lemma C.8. Using the definition of $a_{t+1}$ in Algorithm 3, we have
$\nabla\mathcal{R}_{t}(a_{t+1})-\nabla\mathcal{R}_{t}(a_{t})=\eta_{\rho}\hat{g}_{t}.$
Therefore we have
$\mathbb{E}\left[\hat{g}_{t}^{\top}(a_{t}-a^{*}_{T})-\frac{l_{0}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]$ $=\frac{1}{\eta_{\rho}}\mathbb{E}\left[(\nabla\mathcal{R}_{t}(a_{t+1})-\nabla\mathcal{R}_{t}(a_{t}))^{\top}(a_{t}-a^{*}_{T})-\frac{l_{0}\alpha\eta_{\rho}}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]$
$=\frac{1}{\eta_{\rho}}\mathbb{E}\left[D_{\mathcal{R}_{t}}(a^{*}_{T},a_{t})+D_{\mathcal{R}}(a_{t},a_{t+1})-D_{\mathcal{R}_{t}}(a^{*}_{T},a_{t+1})-\frac{l_{0}\alpha\eta_{\rho}}{4}\|a_{t}-a^{*}_{T}\|^{2}\right].$
$D_{\mathcal{R}}(a,b)$ is the Bregman divergence associated with $\mathcal{R}$, defined by
$D_{\mathcal{R}}(a,b)=\mathcal{R}(a)-\mathcal{R}(b)-\nabla\mathcal{R}(b)^{\top}(a-b).$
Then we can get Equation (22) by using Lemma C.13, C.14, C.15.
$\mathbb{E}\left\{\sum^{T}_{t=1}(\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*}_{T}))\right\}\leq\frac{\nu\log(T)}{\eta_{\rho}}+4d^{2}\eta_{\rho}T+\mathbb{E}\{\sum^{T}_{t=1}b_{t}^{\top}(a_{t}-a^{*}_{T})\}.$ (22)
Lemma C.13 (Lemma 10 in Kumagai (2017)).
Let $\mathcal{R}^{*}_{t}(a)=\sup_{x\in\mathbb{R}^{d}}x^{\top}a-\mathcal{R}_{t}(a)$ denote the Frenchel dual of $\mathcal{R}_{t}$. Then we have
$\sum^{T}_{t=1}\mathbb{E}\left[\hat{g}_{t}^{\top}(a_{t}-a^{*}_{T})-\frac{l_{0}\alpha}{4}\|a_{t}-a^{*}_{T}\|^{2}\right]\leq\frac{1}{\eta_{\rho}}\left(\mathcal{R}(a^{*}_{T})-\mathcal{R}(a_{1})+\mathbb{E}\left[\sum^{T}_{t=1}D_{\mathcal{R}^{*}_{t}}(\nabla\mathcal{R}_{t}(a_{t})-\eta_{\rho}\hat{g}_{t},\nabla\mathcal{R}_{t}(a_{t}))\right]\right).$
Lemma C.14 (Lemma 11 in Kumagai (2017)).
When $\eta_{\rho}\leq\frac{1}{2d}$, we have
$D_{\mathcal{R}^{*}_{t}}(\nabla\mathcal{R}_{t}(a_{t})-\eta_{\rho}\hat{g}_{t},\nabla\mathcal{R}_{t}(a_{t}))\leq 4d^{2}\eta^{2}.$
Lemma C.15 (Lemma 4 in Hazan and Levy (2014)).
$\mathcal{R}(a^{*}_{T})-\mathcal{R}(a_{1})\leq\nu\log(T).$
If we let $C_{0}=(2\nu+4L_{0}\beta+4B_{2}L^{2}+LL_{0}\beta)/\lambda$, we have
$\text{Reg}^{\text{DB}}_{T}$ $\leq\frac{2\nu\log(T)}{\eta_{\rho}}+8d^{2}\eta_{\rho}T+2\mathbb{E}\left\{\sum^{T}_{t=1}b_{t}^{\top}(a_{t}-a^{*}_{T})\right\}+\frac{4L_{0}\beta+4B_{2}L^{2}+LL_{0}\beta}{\lambda\eta_{\rho}}\log T+2LL_{0}R$
[…]
which completes the proof.
C.3.4 Proof for Lemma C.10: Observation Error
Proof  (Proof for Lemma C.10) In the following, we will analyze the expected cumulative sum of the bias $b_{t}$ on the total regret, which is $\mathbb{E}\left\{\sum^{T}_{t=1}b^{\top}_{t}(a_{t}-a^{*}_{T})\right\}$. By Cauchy-Schwarz inequality, $b^{\top}_{t}(a_{t}-a^{*}_{T})\leq\|b_{t}\|_{2}\|(a_{t}-a^{*}_{T})\|_{2}$. We can bound the $\|b_{t}\|_{2}$ by the following.
$b^{\top}_{t}b_{t}$ $=d^{2}\left(P_{t}(a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t})-\hat{P}_{t}(a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t})\right)^{2}u^{\top}_{t}\nabla^{2}\mathcal{R}_{t}(a_{t})u_{t}$
$\leq\left(P_{t}(a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t})-\hat{P}_{t}(a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t})\right)^{2}d^{2}\lambda_{\max}(\nabla^{2}\mathcal{R}_{t}(a_{t})).$
We get the first equality by Lemma C.8. We get the inequality by using the fact $\|u_{t}\|_{2}=1$. If we let $a^{\prime}_{t}=a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t}$, then we have
$|P_{t}(a^{\prime}_{t})-\hat{P}_{t}(a^{\prime}_{t})|$ $=|\sigma(f(a_{t})-f(a^{\prime}_{t}))-\sigma(f(a_{t})-f(a^{\prime}_{t})+c_{t}(a_{t},a^{\prime}_{t}))|$
$\leq\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right).$
We get the first equality by using the definition of $P_{t}(a_{t})$ and $\hat{P}_{t}(a_{t})$. We get the first inequality by using the Lipschitz property of $\sigma$ (see Assumption C.3) Therefore, we have
$\|b_{t}\|_{2}\leq d\sqrt{\lambda_{\max}(\nabla^{2}\mathcal{R}_{t}(a_{t}))}\min\left\{2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right\}.$
If we use $\lambda^{*}_{R}:=\sup_{a\in\mathcal{A}}\lambda_{\max}\left(\nabla^{2}\mathcal{R}(a)\right)$, then we have
$\lambda_{\max}(\nabla^{2}\mathcal{R}_{t}(a_{t}))$ $=\lambda_{\max}(\nabla^{2}\mathcal{R}(a_{t}))+\lambda\eta_{\rho}t+2\mu\leq\lambda^{*}_{R}+2\mu+\lambda\eta_{\rho}t.$
Therefore we have
$\mathbb{E}\left\{\sum^{T}_{t=1}b^{\top}_{t}(a_{t}-a^{*}_{T})\right\}\leq\mathbb{E}\left\{\sum^{T}_{t=1}\|b_{t}\|_{2}\|a_{t}-a^{*}_{T}\|_{2}\right\}$
$\leq\mathbb{E}\left\{\sum^{T}_{t=1}(d\sqrt{\lambda_{\max}(\nabla^{2}\mathcal{R}_{t}(a_{t}))}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right))\|a_{t}-a^{*}_{T}\|_{2}\right\}$
$\leq\mathbb{E}\left\{\sum^{T}_{t=1}(d\sqrt{\lambda^{*}_{R}+2\mu+\lambda\eta_{\rho}t}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right))\|a_{t}-a^{*}_{T}\|_{2}\right\}$
$\leq\mathbb{E}\left\{\sum^{T}_{t=1}(d\sqrt{\lambda^{*}_{R}+2\mu}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\|a_{t}-a^{*}_{T}\|_{2}\right\}+\mathbb{E}\left\{\sum^{T}_{t=1}(d\sqrt{\lambda\eta_{\rho}t}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\|a_{t}-a^{*}_{T}\|_{2}\right\}$
$\leq 2RdL_{\sigma}\sqrt{\lambda^{*}_{R}+2\mu}C+d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\|a_{t}-a^{*}_{T}\|_{2}\right\}$ (23)
$\leq 2RdL_{\sigma}\sqrt{\lambda^{*}_{R}+2\mu}C+2d\sqrt{\lambda\eta_{\rho}T}+d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\|a_{t}-a^{*}\|_{2}\right\}$ (24)
$\leq 2RdL_{\sigma}\sqrt{\lambda^{*}_{R}+2\mu}C+2d\sqrt{\lambda\eta_{\rho}T}+d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\min\left(2R,\sqrt{\frac{2}{\alpha}\left(f(a_{t})-f(a^{*})\right)}\right)\right\}.$
We get Equation (23) by using the fact that $\|a_{t}-a^{*}_{T}\|_{2}\leq 2R$, $\sum^{T}_{t=1}|c_{t}(a_{t},a^{\prime}_{t})|\leq C$. We get Equation (24) the definition of $a^{*}_{T}$, and the fact $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$. Using the $\alpha$-strong convexity of $f$ (see Assumption C.2), we have $\|a_{t}-a_{*}\|_{2}\leq\min\left(2R,\sqrt{\frac{2}{\alpha}\left(f(a_{t})-f(a^{*})\right)}\right)$ we get the last inequality, which completes the proof.
C.3.5 Proof for Lemma C.11
Proof  (Proof for Lemma C.11). Using Abel’s Summation Equation (Lemma C.16), we have
$\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}\leq\sqrt{T}\left(\sum^{T}_{t=1}m_{t}n_{t}\right)$
Since we have $\sum^{T}_{t=1}n^{2}_{t}\leq C^{\prime}T^{\frac{1}{2}+\alpha}$ and $0\leq n_{t}\leq K,\forall t$, we have $\sum^{T}_{t=1}n_{t}\leq\sqrt{C^{\prime}}T^{\frac{3}{4}+\frac{\alpha}{2}},\forall t$. Moreover, for $t\geq 1$, $t^{-1+\rho}-(t+1)^{-1+\rho}\leq t^{\rho-\frac{1}{2}}$ holds for all $t\geq 1$. This is because
$f(t)=t^{-1+\rho}\left((1+\frac{1}{t})^{-1+\rho}+\frac{1}{t}-1\right).$
is decreasing over $t\geq 1$ and $\lim_{t\rightarrow\infty}f(t)=0$. Consequently, applying Abel’s Summation Equation again, we have
$\sum^{T}_{t=1}m_{t}n_{t}$ $=\left(\sum^{T}_{t=1}n_{t}\right)m_{T}+\sum^{T-1}_{t=1}\left(\sum^{t}_{i=1}n_{i}\right)(m_{t}-m_{t+1})$
$\leq\sqrt{C^{\prime}}c_{k}T^{\frac{3}{4}+\frac{\alpha}{2}}T^{-1+\rho}+\sqrt{C^{\prime}}c_{k}\sum^{T-1}_{t=1}\left(\sum^{t}_{i=1}n_{i}\right)\left(t^{-1+\rho}-(t+1)^{-1+\rho}\right)$
$\leq\sqrt{C^{\prime}}c_{k}T^{\frac{3}{4}+\frac{\alpha}{2}}T^{-1+\rho}+\sqrt{C^{\prime}}c_{k}\sum^{T-1}_{t=1}\left(\sum^{t}_{i=1}n_{i}\right)t^{-\frac{1}{2}+\rho}$
$\leq\sqrt{C^{\prime}}c_{k}T^{\frac{3}{4}+\frac{\alpha}{2}}T^{-1+\rho}+\sqrt{C^{\prime}}c_{k}\sum^{T-1}_{t=1}t^{\frac{3}{4}+\frac{\alpha}{2}}t^{-\frac{1}{2}+\rho}$
$\leq 5\sqrt{C^{\prime}}c_{k}T^{-\frac{1}{4}+\frac{\alpha}{2}+\rho}.$
Since $0\leq n_{t}\leq K$, $\sum^{t}_{i=1}n_{i}$ is increasing and the increasing rate is upper bound by $t$. Together with the constraint $\sum^{T}_{t=1}n_{t}\leq\sqrt{C^{\prime}}T^{\frac{3}{4}+\frac{\alpha}{2}}$, $\sum^{T-1}_{t=1}\left(\sum^{t}_{i=1}n_{i}\right)t^{-\frac{1}{2}+\rho}$ is optimized when $\sum^{t}_{i=1}n_{i}$ increasing in the speed of $\sqrt{C^{\prime}}t^{\frac{3}{4}+\frac{\alpha}{2}}$. Because of this, the second last inequality holds.
Lemma C.16.
(Abel’s Summation Equation (Williams, 1963)). For any numbers $a_{k}$, $b_{k}$, we have
$\sum^{n}_{k=1}a_{k}b_{k}=\left(\sum^{n}_{k=1}b_{k}\right)a_{n}+\sum^{n-1}_{k=1}\left(\sum^{k}_{i=1}b_{i}\right)(a_{k}-a_{k+1}).$
C.4 Proof for Proposition 2: Efficiency-Robustness Tradeoff in NC-SMD
Theorem (Full Statement of Proposition 2).
Let the number of rounds $T$ be fixed and $T>\sqrt{2LL_{0}R}$. Under Assumptions C.2, C.3, and C.4, by choosing $\lambda\leq l_{0}\alpha/2$, $\mu\geq(L^{3}_{0}L_{2}/\lambda)^{2}$, and setting the learning parameter $\eta_{\rho}=\frac{\sqrt{\log T}}{2d}T^{-\alpha},\alpha\in[0.5,1]$, regret occurred by Algorithm 3 satisfies
$\text{Reg}_{T}\leq\tilde{O}\left(dT^{\alpha}+\sqrt{d}T^{\frac{1}{2}(1-\alpha)}C+dC\right)$
for any corruption with budget $C$.
Proof  (Proof for Proposition 2). From Lemma C.10, we know
$\mathbb{E}\left\{\sum^{T}_{t=1}b^{\top}_{t}(a_{t}-a^{*}_{T})\right\}$ $\leq 2RdL_{\sigma}\sqrt{\lambda^{*}_{R}+2\mu}C+d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\|a_{t}-a^{*}_{T}\|_{2}\right\}$
$\leq 2RdL_{\sigma}\sqrt{\lambda^{*}_{R}+2\mu}C+2Rd\sqrt{\lambda\eta_{\rho}}L_{\sigma}\sum^{T}_{t=1}\sqrt{t}|c_{t}(a_{t},a^{\prime}_{t})|$
$\leq 2RdL_{\sigma}\sqrt{\lambda^{*}_{R}+2\mu}C+2Rd\sqrt{\lambda\eta_{\rho}T}L_{\sigma}C.$
We get the second inequality by using the fact that the diameter of the action space of $R$. We get the last inequality by using Abel Summation Equation (see Lemma C.16). Therefore, by Lemma C.9, we have
$\text{Reg}_{T}\leq\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2LL_{0}R+4RdL_{\sigma}\sqrt{\lambda^{*}_{R}+2\mu}C+4Rd\sqrt{\lambda\eta_{\rho}T}L_{\sigma}C.$
Choosing $\eta_{\rho}=\frac{\sqrt{\log T}}{2d}T^{-\alpha},\alpha\in[0.5,1]$, we have
$\text{Reg}_{T}$ $\leq 2dC_{0}\sqrt{\log T}T^{\alpha}+4d\sqrt{\log T}T^{1-\alpha}+4RdL_{\sigma}\sqrt{\lambda^{*}_{R}+2\mu}C+4R\sqrt{d}L_{\sigma}(\log T)^{0.25}T^{\frac{1}{2}(1-\alpha)}C+2LL_{0}R$
$\leq O\left(d\sqrt{\log T}T^{\alpha}+\sqrt{d}(\log T)^{\frac{1}{4}}T^{\frac{1}{2}(1-\alpha)}C+dC\right),$
which completes the proof.
D Additional Experiments
In this section, we list all the experiment details. At the beginning, we introduce the regret order fitting method, different types of corruption that we simulate in the experiments, and baseline algorithms for comparison.
Fitted Order of Regret. We introduce the methodology that we use to compute the order of $\text{Reg}_{T}$ in terms of the number of iteration, $T$. To fit the order of the regret, we first convert it into $\log$ scale. Then we input the last $1\%$ of data, run linear regression, and use ordinary least squares to estimate the slope of the line, which is the fitted order of $\text{Reg}_{T}$.
Corruption. In the experiments, we consider two types of corruption: corruption induced by $\rho$-imperfect user and arbitrary corruption. For corruption induced by $\rho$-imperfect user, we simulate $c_{t}(a_{t},a^{\prime}_{t})$ according to Definition A.1. We set $\lambda=2$ and $C_{0}=0.1(\max_{a\in\mathcal{A}}\mu(a)-\min_{a\in\mathcal{A}}\mu(a))$. Moreover, we add $c_{t}(a_{t},a^{\prime}_{t})$ to the utility difference of $a_{t}$ and $a^{\prime}_{t}$ adversarially. In particular, if $\mu(a_{t})>\mu(a^{\prime}_{t})$, then ^$\hat{\mathbb{P}}(a_{t}\succ a^{\prime}_{t})=\sigma(\mu(a_{t})-\mu(a^{\prime}_{t})-c_{t}(a_{t},a^{\prime}_{t}))$, then vice versa. For arbitrary corruption, we force the user to submit her least preferred item each round to the algorithm over the first $C$ rounds.
Baseline Algorithms. We consider three baseline algorithms for comparison, Doubler (Ailon et al., 2014), Sparring (Ailon et al., 2014), and Versatile-DB (Saha and Gaillard, 2022).
Doubler is the first approach that transforms a dueling bandit problem into a standard multi-armed bandit (MAB) problem. It operates in epochs of exponentially increasing length: in each epoch, the left arm is sampled from a fixed distribution, and the right arm is selected using a MAB algorithm to minimize regret against the left arm. The feedback received by the MAB algorithm is the number of wins the right arm achieves compared to the left arm. Under linear link assumption, Doubler has been proven to experience regret as the same order as underlying MAB algorithm. For continuous action space and general concave utility, we choose Bandit Gradient Descent (BGD, (Flaxman et al., 2004)), with regret $O(T^{3/4})$, as the underlying MAB algorithm.
Sparring initializes two MAB instances and lets them compete against each other. It is a heuristic improvement over Doubler. Although it does not come with a regret upper bound guarantee, it is reported to enjoy better performance compared to Doubler (Ailon et al., 2014). We also choose BGD as the underlying MAB algorithm.
Versatile-DB applies novel reduction approach which converts dueling bandit to MAB. Essentially, it designs simplex over $K$ arms using the idea of follow-the-regularized-leader and updates the simplex using information of whether the arm wins. It has been proved that Versatile-DB is robust to corruption defined as the number of flips of duels and Versatile-DB enjoys regret linear in $C$ (Saha and Gaillard, 2022).
D.1 Experiments on Synthetic Data
D.1.1 DBGD Lower Bound Simulation
Experiment Setup. To validate Corollary 1, we construct the following problem instance. Consider the linear utility function $\mu(a):=\theta^{\top}a$. Choose $d=2,\theta=[\frac{1}{2},\frac{1}{2}],T=10^{5}$. Let action set $\mathcal{A}:=\{(a_{1},a_{2}):a_{1}\geq 0,a_{2}\geq 0,\frac{1}{2}a_{1}+a_{2}-\frac{1}{4}\leq 0\}$. We repeat the experiments for $50$ times, each time with a different seed.
Results and Discussion. Each grey-colored dotted line in Figure 2 indicate the cumulative regret for each simulation. The blue line is the average cumulative regret over 50 repeats. The error bar indicates $\pm$ one standard deviation of the regret. The fitted order over the last $1\%$ of the average cumulative regret is 0.76, which is close $0.75$ and supports Corollary 1.
D.1.2 Regret Upper Bound in Theorem 1
Experiment Setup. To validate regret upper bound in Theorem 1, we consider a strongly concave utility $\mu_{\theta}(a):=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$, a logistic link function $\sigma(x)=\frac{1}{1+\exp(-x)}$. We choose $d=5$, and $T=10^{5}$. Our action space $\mathcal{A}$ is a $d$-dimensional ball with radius $R=10$. The preference parameter $\theta$ is randomly sampled from the surface of $\mathcal{A}$. In our problem setting, the optimal action $a^{*}$ is $\theta$ and $\mu_{\theta}(a^{*})=50$. We simulate corruption induced by $\rho$-imperfect user and conduct experiments for $\rho\in[0.5,1.0]$. For each value of $\rho$, we repeat the experiments for $5$ times, each under a different seed. We set the learning rate $\eta_{\rho}$ according to Proposition C.1.
Figure D.1: Performance under Known C
Results and Discussion. Figure D.1 shows that when $\rho$ is known, NC-SMD can tolerate linear $\rho$-imperfect user corruption, which confirms theoretical analysis in Proposition C.1 and implies that lower bound in Theorem 1 is tight.
D.1.3 Robustness under Corruption induced by Imperfect User
In Figure D.2, we simulated the performance of DBGD, NC-SMD, Sparring, Doubler in the same experiment setting as in Section 6.1 expect using corruption induced by $\rho$-Imperfect User. We notice that the experiment results also align with theoretical prediction (Theorem 2 and Proposition 2), which predicts when $\alpha=0.25$ for DBGD and $\alpha=0.5$ for NC-SMD, they can tolerate $O(T^{0.75})$ agnostic corruption.
Figure D.2: Robustness under Corruption induced by $\rho$-Imperfect User
D.1.4 Efficiency-Robustness Tradeoff in Theorem 2
To validate efficiency-robustness trade-off in Theorem 2 and Proposition 2, we adopt the same experiment setup as in Section 6.1 except different $\alpha$ values. In the first row of Figure D.3, we consider $\alpha=0.05$ for DBGD and $\alpha=0.9$ for NC-SMD. We consider $\rho\in[0.5,0.95]$. According to the theoretical prediction, both DBGD and NC-SMD can tolerate at most $O(T^{0.95})$ agnostic corruption, which aligns with experiment results. In the second row of Figure D.3, we consider $\alpha=0.1$ for DBGD and $\alpha=0.8$ for NC-SMD. According to the theoretical prediction, both DBGD and NC-SMD can tolerate at most $O(T^{0.9})$ agnostic corruption. From the figure, we can see they have smaller fitted order of regret when $\rho=0.5$ while at a cost of tolerating smaller magnitude of agnostic corruption (it has linear regret when $\rho=0.95$), which reveals intrinsic tradeoff between efficiency and robustness.
Figure D.3: Efficiency-Robustness Tradeoff
D.2 Experiments on Spotify Recommendation Data
Evaluation Setup. We evaluate our approach on Spotify recommendation data (Spotify, 2020). The objective is to recommend songs to incoming users. This dataset includes $17\times 10^{4}$ songs ($|\mathcal{A}|=17\times 10^{4}$), each is described by 15 distinct features (d = 15).
User Profile Identification. In order to model different types of user, we first use Standard Scaler to standardize the dataset. Then we use Kmeans to conduct clustering (Pedregosa et al., 2011).
User Utility. We use the average of the song embedding within each group as the preference vector. Our aim is to recommend songs to users that have the highest cosine similarity (utility function) with their preference vectors. We rescale the cosine similarity to $[-100,100]$.
D.2.1 Performance of Versatile-DB on Spotify Recommendation Data
To test the performance of Versatile-DB, we randomly sample $K=2000$ songs from the Spotify Recommendation Dataset. We identify 2 different user types by clustering the dataset into 2 groups. We use the average of the song embedding within each group as the preference vector and compute the user utility for each recommendation. We consider corruption induced by $\rho$-imperfect user for $\rho\in[0.5,0.75]$. We set $\alpha=0.25$ for DBGD and run both algorithms for $T=100$ iterations. This is because Versatile-DB is extremely computationally intensive and running $K=2000,T=100$ takes around 20 CPU hours. For each user type and for each value of $\rho$, we repeat the experiment over 5 times, each under a different seed. From Figure D.4, we notice that DBGD outperforms Versatile-DB in all $\rho$, which highlights DBGD’s robustness and applicability in real-world data.
Figure D.4: Performance Comparison between DBGD and Versatile-DB